### v1

# On the controller form for linear hyperbolic MIMO systems with dynamic boundary conditions

Abstract
This contribution develops an algebraic approach to obtain a controller form for a class of linear hyperbolic MIMO systems, bidirectionally coupled with a linear ODE system at the unactuated boundary. After a short summary of established controller forms for SISO and MIMO ODE as well as SISO hyperbolic PDE systems, it is shown that the direct approach to state a controller form fails already for a very simple MIMO example. Next, a generalised hyperbolic controller form with different variants is proposed and a new flatnesss-based scheme to compute said form is presented. Therein, the system is treated in an algebraic setting where generalised polynomials with real exponents are used to describe the predictions and delays in the system. The proposed algorithm is then applied to the motivating example.
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
I Introduction
From the onset of state space based approaches, controller (canonical) forms have been a vital asset for systematic controller designs. As the controller forms for linear SISO and MIMO ODE systems, cf. [1, 2] as well for SISO hyperbolic PDE systems [3] are long established and understood, only comparatively recent works [4, 5, 6] discuss hyperbolic systems with boundaries governed by ODEs for the SISO case which this contribution extends to the MIMO case.
Following some preliminaries, where we briefly restate established controller forms and demonstrate the problems that arise if a hyperbolic controller form is to be introduced even for the most simple MIMO system, a MIMO hyperbolic controller form (HCF) is given in Section III. The main part of the contribution in Section IV then discusses how a HCF can be obtained for a specific class of systems whereupon Section V tracks the procedure for an example system.
II Preliminaries
II-A Notation
The variables $z$ and $t$ denote space and time. The first and second derivative of $x(t)$ is denoted by $\dot{x}(t)$ and $\ddot{x}(t)$, while the $n$-th derivative is given by $x^{(n)}(t)$ and the partial derivative of $x(z,t)$ w.r.t. $z$. by $\partial_{z}x(z,t)$.
To treat delay differential equations (DDEs) in an algebraic setting, we identify the operators therein with their Fourier-Laplace transforms and treat the application of the transformed operators as convolutions. Thus, we write $\mathcal{L}\{\partial_{t}\}=s$ for differentials, while the shifts $(\sigma^{\tau}x)(t)=x(t+\tau)$ with $\tau\in\mathbb{R}$ are denoted as $\mathcal{L}\{\sigma^{\tau}\}(s)=\exp(s\tau)$.
The shorthand $L^{2}([a,b])$ is used for the space of square integrable functions on $[a,b]$, while $H^{n}([a,b])$ stands for the space of functions whose $n$-th derivative belongs to $L^{2}([a,b])$ and $\mathcal{E}^{\prime}$ for Schwartz-Distributions with compact support. With $\operatorname{V}_{f}([a,b])$ denoting the total variation of a function $f$ on the interval $[a,b]$, $BV([a,b])$ contains functions of bounded variation while members of its subset $BV^{\star}([a,b])$ satisfy $\lim_{\epsilon\to 0}\operatorname{V}_{f}([b,b+\epsilon))=0$.
Furthermore, $\mathbb{R}[\partial_{t}]$ is the space of polynomials in $\partial_{t}$ with coefficients in $\mathbb{R}$ and $\mathbb{R}(\partial_{t})$ consists of rational expressions in $\partial_{t}$. Regarding a polynomial matrix $X(\partial_{t})\in\mathbb{R}^{n\times n}[\partial_{t}]$, we denote the leading column coefficient matrix (LCCM) w.r.t. $\partial_{t}$ by $\operatorname{lccm}_{\partial_{t}}{\mathbf{X}(\partial_{t})}$ and say that $\mathbf{X}(\partial_{t})$ is column reduced w.r.t. $\partial_{t}$ if $\operatorname{lccm}_{\partial_{t}}{\mathbf{X}(\partial_{t})}$ has full generic rank. Finally, we call $\mathbf{X}(\partial_{t})$ unimodular if its determinant is independent of $\partial_{t}$.
II-B Established Controller Forms
II-B1 ODE SISO systems
For this class, [1] defines the controller canonical form
$\dot{\eta}^{\mathrm{O}}_{1}(t)$ $=\eta^{\mathrm{O}}_{2}(t)$ (1a)
$\;\;\vdots$
$\dot{\eta}^{\mathrm{O}}_{n-1}(t)$ $=\eta^{\mathrm{O}}_{n}(t)$ (1b)
$\dot{\eta}^{\mathrm{O}}_{n}(t)$ $=-\bm{a}^{\intercal}\bm{\eta}^{\mathrm{O}}(t)+u(t)$ (1c)
with state $\bm{\eta}^{\mathrm{O}}(t)=(\eta^{\mathrm{O}}_{1}(t),\dotsc,\eta^{\mathrm{O}}_{n}(t))^{\intercal}\in\mathbb{R}^{n}$, coefficients $\bm{a}\in\mathbb{R}^{n}$ and input $u(t)\in\mathbb{R}$ that exists for every controllable system. Note that the “output” $\eta^{\mathrm{O}}_{1}(t)$ of the integrator chain is a flat output of the system.
II-B2 ODE MIMO systems
Extending the SISO framework to systems with inputs $\bm{u}(t)\in\mathbb{R}^{m}$, [2] states a controller form [^1] As these forms are not unique in the MIMO case, the ancillary adjective canonical will be omitted from now on. Moreover, note that (2c) explicitly allows for all inputs to occur in the “input” of the subsystems integrator chain with linear factors $\neq 1$ as long as $\mathbf{B}=(\bm{b}_{1},\dotsc,\bm{b}_{m})^{\intercal}$ has full rank. , with subsystems
$\dot{\eta}^{\mathrm{O}}_{i,1}(t)$ $=\eta^{\mathrm{O}}_{i,2}(t)$ (2a)
$\;\;\vdots$
$\dot{\eta}^{\mathrm{O}}_{i,\nu_{i}-1}(t)$ $=\eta^{\mathrm{O}}_{i,\nu_{i}}(t)$ (2b)
$\dot{\eta}^{\mathrm{O}}_{i,\nu_{i}}(t)$ $=-\sum_{k=1}^{m}\bm{a}_{k}^{\intercal}\bm{\eta}^{\mathrm{O}}_{k}(t)+\bm{b}_{i}^{\intercal}\bm{u}(t)$ (2c)
for $i\in\{1,\dotsc,{m}\}$, with the Kronecker index $\nu_{i}$ and the state vector $\bm{\eta}^{\mathrm{O}}_{i}(t)=(\eta^{\mathrm{O}}_{i,1}(t),\dotsc,\eta^{\mathrm{O}}_{i,\nu_{i}}(t))^{\intercal}\in\mathbb{R}^{\nu_{i}}$ of the $i$-th subsystem. Similarly to the SISO case, the MIMO controller form state $\bm{\eta}^{\mathrm{O}}(t)\in\mathbb{R}^{\nu_{\Sigma}}$ with $\nu_{\Sigma}\coloneq\sum_{i=1}^{m}\nu_{i}$ and input $\bm{u}(t)$ can be easily parametrised by the tuple $(\eta^{\mathrm{O}}_{1,1}(t),\dotsc,\eta^{\mathrm{O}}_{m,1}(t))$.
II-B3 hyperbolic SISO systems
Initially, the work [3] introduced a controller form for boundary controlled hyperbolic systems that takes the form of decoupled transport systems. These results were then extended to dynamic boundary conditions in [4, 5], yielding the controller form
$\dot{\eta}^{\mathrm{O}}_{1}(t)$ $=\eta^{\mathrm{O}}_{2}(t)$ (3a)
$\;\;\vdots$
$\dot{\eta}^{\mathrm{O}}_{n-1}(t)$ $=\eta^{\mathrm{O}}_{n}(t)$ (3b)
$\dot{\eta}^{\mathrm{O}}_{n}(t)$ $=\eta^{\mathrm{P}}(\check{\tau},t)$ (3c)
[…]
$\eta^{\mathrm{P}}(\hat{\tau},t)$ $=-\operatorname{\mathcal{S}}_{a^{\star}}\eta^{\mathrm{P}}(\cdot,t)-\bm{a}^{\intercal}\eta^{\mathrm{O}}(t)+u(t).$ (3e)
[…]
II-C Benchmark System
Consider a system of two vibrating strings
$\partial_{t}^{2}x_{1}(z,t)-\partial_{z}^{2}x_{1}(z,t)$ $=0\quad z\in(-\alpha,0)$ (4a)
$\partial_{t}^{2}x_{2}(z,t)-\partial_{z}^{2}x_{2}(z,t)$ $=0\quad z\in(0,\beta),$ (4b)
coupled with a point mass at $z=0$ and thus subject to the dynamic boundary conditions
$x_{1}(0,t)=x_{2}(0,t)$ $=\xi(t)$ (5a)
$m\ddot{\xi}(t)$ $=-\partial_{z}x_{1}(0,t)+\partial_{z}x_{2}(0,t),$ (5b)
where $\xi(t)$ denotes the mass position as well as
$\partial_{z}x_{1}(-\alpha,t)$ $=-u_{1}(t)$ (5c)
$\partial_{z}x_{2}(\beta,t)$ $=u_{2}(t),$ (5d)
with state [^2] In detail, $\bm{x}(\cdot,t)\!=\!(\xi(t),\dot{\xi}(t),x_{1}(\cdot,t),\partial_{t}x_{1}(\cdot,t),x_{2}(\cdot,t),\partial_{t}x_{2}(\cdot,t))^{\intercal}$ and $\mathcal{X}=\{\mathbb{R}^{2}\times L^{2}([-\alpha,0])\times H^{1}([-\alpha,0])\times L^{2}([0,\beta])\times H^{1}([0,\beta])\;|\;\xi(t)=x_{1}(0,t)=x_{2}(0,t)\}$. $\bm{x}(\cdot,t)\in\mathcal{X}$ as initially discussed in [7]. By using the deflection and gradient at $z=0$ as Cauchy data, D’Alemberts’s formula yields the solutions
$x_{i}(z,t)\!=\!\frac{1}{2}\!\left(x_{i}(0,t-z)\!+\!x_{i}(0,t+z)\right)\!+\!\frac{1}{2}\!\int\limits_{t-z}^{t+z}\partial_{z}x_{i}(0,\tau)\,\mathrm{d}\tau$ (6)
for $i=1,2$ (c.f. [7, 4]), which reads
$x_{i}(z,t)=C_{z}x_{i}(0,t)+\frac{1}{s}S_{z}\partial_{z}x_{i}(0,t)$ (7)
in the algebraic setting, where $C_{z}\coloneqq\cosh(\partial_{t}z)=\frac{1}{2}(\sigma^{z}+\sigma^{-z})$ and $S_{z}\coloneqq\sinh(\partial_{t}z)=\frac{1}{2}(\sigma^{z}-\sigma^{-z})$.
II-D Parametrisation
Starting at the interior mass, by using the tuple
$\begin{pmatrix}{y}_{1}(t),{y}_{2}(t)\end{pmatrix}^{\intercal}=\begin{pmatrix}\xi(t),\partial_{z}x_{1}(0,t)+\partial_{z}x_{2}(0,t)\end{pmatrix}^{\intercal}$ (8)
with $y_{1}\in H^{2}(\mathbb{R})$ and $y_{2}\in L^{2}(\mathbb{R})$ we directly obtain $\xi(t)=x_{1}(0,t)=x_{2}(0,t)={y}_{1}(t)$ as well as
$\partial_{z}x_{1}(0,t)$ $=\frac{1}{2}\left({y}_{2}(t)-m\partial_{t}^{2}{y}_{1}(t)\right)$ (9a)
$\partial_{z}x_{2}(0,t)$ $=\frac{1}{2}\left({y}_{2}(t)+m\partial_{t}^{2}{y}_{1}(t)\right)$ (9b)
from (5b). With the Cauchy data known, (7) then yields the profiles, from which the input parameterisation
$\begin{pmatrix}u_{1}(t)\\ u_{2}(t)\\ \end{pmatrix}=\begin{pmatrix}\frac{m}{2}\partial_{t}^{2}C_{\alpha}+\partial_{t}S_{\alpha}&-\frac{1}{2}C_{\alpha}\\ \frac{m}{2}\partial_{t}^{2}C_{\beta}+\partial_{t}S_{\beta}&\frac{1}{2}C_{\beta}\end{pmatrix}\begin{pmatrix}y_{1}(t)\\ y_{2}(t)\\ \end{pmatrix}$ (10)
is obtained via (5c) and (5d) showing that (8) is a flat output of the system (4)-(5).
II-E Stating a controller form
As known from, e.g [4], the problem to be solved for stating a controller form is to find a particular realisation of the input-output relation (10). A possible way to do so is to solve (10) for the highest derivatives of the largest predictions of $y_{1}(t)$ as well as $y_{2}(t)$ and to check for which initial conditions the resulting DDE system is well posed.
However, taking the coefficients of the highest derivatives of the largest predictions in each column on the right hand-side of (10), the resulting leading column coefficient matrix
$\operatorname{lccm}_{\rho,s}{\mathbf{H}}=\frac{1}{4}\begin{pmatrix}0&0\\ m&1\\ \end{pmatrix}$ (11)
does not have full rank. This implies that (10) cannot directly be written as an explicit DDE system thus negating our approach. Therefore, more steps have to be taken for MIMO systems, which will be discussed in Section IV after we give a definition of the desired controller form for this case.
III A hyperbolic controller form
In this section we define a general controller form for hyperbolic ODE-PDE MIMO systems and state several specific variants of the latter which may be helpful in different settings.
III-A Definition
By directly applying the extension of the SISO controller form by a transport system from (3) to the MIMO controller form (2), for $i\in\{1,\dotsc,m\}$, we obtain the subsystems
$\dot{\eta}^{\mathrm{O}}_{i,1}(t)$ $=\eta^{\mathrm{O}}_{i,2}(t)$ (12a)
$\;\;\vdots$
$\dot{\eta}^{\mathrm{O}}_{i,\nu_{i}-1}(t)$ $=\eta^{\mathrm{O}}_{i,\nu_{i}}(t)$ (12b)
$\dot{\eta}^{\mathrm{O}}_{i,\nu_{i}}(t)$ $=\eta^{\mathrm{P}}_{i}(\check{\tau}_{i},t)$ (12c)
$\partial_{t}\eta^{\mathrm{P}}_{i}(\tau,t)$ $=\partial_{\tau}\eta^{\mathrm{P}}_{i}(\tau,t),\quad\tau\in[\check{\tau}_{i},\hat{\tau}_{i})$ (12d)
$\begin{split}\eta^{\mathrm{P}}_{i}(\hat{\tau}_{i},t)&=-\sum\limits_{j=1}^{m}\operatorname{\mathcal{S}}_{a^{\star}_{i,j}}\eta^{\mathrm{P}}_{j}(\cdot,t)+\bm{a}_{i,j}^{\intercal}\eta^{\mathrm{O}}_{j}(t)\\ &\qquad\qquad+\operatorname{\mathcal{S}}_{b_{i,j}^{\star}}\!*\,u_{j}(t),\end{split}$ (12e)
with state $\bm{\eta}(\cdot,t)\in\mathbb{R}^{\nu_{\Sigma}}\times\prod_{i=1}^{m}L^{2}([\check{\tau}_{i},\hat{\tau}_{i}])$. Herein, the convolutions $\operatorname{\mathcal{S}}_{b_{i,j}^{\star}}\!*\,u_{j}(t)=\int_{\underline{\tau\!}}^{\bar{\tau}}u_{j}(t-\tau)\,\mathrm{d}b_{i,j}^{\star}(\tau)$ with the $b_{i,j}^{\star}$ to be further specified form the input map $\mathbf{B}^{\star}=[\operatorname{\mathcal{S}}_{b^{\star}_{i,j}}]$ which is invertible w.r.t. convolution.
In (12), each block consist of a transport system whose output feeds into an integrator chain. However, the input of each block’s transport system now depends on the system input as well as the PDE and ODE states of all blocks. In the following, we will derive several special cases of the form (12) by specifying the functions $b_{i,j}^{\star}$ that define $\mathbf{B}^{\star}$.
III-B A classic controller form
To recover a classic form without any input derivatives, predictions or delays, we restrain the functions to $b_{i,j}^{\star}(t)=b_{i,j}H(t)$ with the Heavyside function $H(t)$, resulting in
$\sum\limits_{j=1}^{m}\operatorname{\mathcal{S}}_{b^{\star}_{i,j}}\!*\,{u_{j}}(t)=\bm{b}_{i}^{\intercal}\bm{u}(t)$ (13)
for $i\in\{1,\dotsc,{m}\}$ after evaluating the convolutions.
III-C Generalised controller forms
III-C1 With input derivatives
To allow for input derivatives to occur in (12e), we set $b_{i,j}^{\star}(t)=\sum_{k=0}^{D_{i,j}}b_{k}H^{(k)}(t)$, yielding
$\sum\limits_{j=1}^{m}\operatorname{\mathcal{S}}_{b^{\star}_{i,j}}\!*\,{u_{j}}(t)=\sum\limits_{j=1}^{m}\sum_{k=0}^{D_{i,j}}b_{i,j,k}u_{j}^{(k)}(t)$ (14)
where the $D_{i,j}$ denote the highest derivate order of $u_{j}(t)$ in the equations of subsystem $i$. Owing to the derivatives of the $u_{j}\in L^{2}(\mathbb{R}^{+})$, we propose the term discontinuous controller form for the subtype (14).
III-C2 With input predictions
For this case, we have $b_{i,j}^{\star}\in BV([-\bar{\tau}_{i,j},0])$, where the $\bar{\tau}_{i,j}$ denote the biggest predictions with which the input $u_{j}(t)$ appears in the subsystem $i$. As such a system depends on future inputs, we propose the term non-causal controller form for this case.
III-C3 With input predictions and delays
For a more general case, we allow $b_{i,j}^{\star}\in BV([-\bar{\tau}_{i,j},-\underline{\tau\!}_{i,j}])$, where the $\underline{\tau\!}_{i,j}$ denote the biggest delays of input $u_{j}(t)$. As this form only yields a sensible state if the system is defined for $t\in\mathbb{R}$ instead of $\mathbb{R}^{+}$, we propose the term quasi controller form.
IV Computation of the controller form
This section contains the main result of the contribution, in which we present an algorithm to compute the controller form for a specific class of systems.
IV-A System Class
We discuss the linear first order hyperbolic MIMO system [^3] Note that the system from Section II-C can easily be transformed into this more general form which also allows for unsymmetrical predictions and delays in contrast to the more compact second order formulation.
$\dot{\bm{\xi}}(t)$ $=\bm{F}\bm{\xi}(t)+\bm{B}\bm{x}^{-}(0,t)$ (15a)
$\partial_{t}\bm{x}(z,t)$ $=\bm{\Lambda}\partial_{z}\bm{x}(z,t),\quad z\in(0,1)$ (15b)
$\bm{x}^{+}(0,t)$ $=\bm{Q}_{0}\bm{x}^{-}(0,t)+\bm{C}\bm{\xi}(t)$ (15c)
$\bm{x}^{-}(1,t)$ $=\bm{Q}_{1}\bm{x}^{+}(1,t)+\bm{u}(t)$ (15d)
as depicted in Figure 1, where $\bm{\xi}(t)\in\mathbb{R}^{n}$, $\bm{x}^{-}(\cdot,t)\in L^{2}([0,1])^{{n^{-}}}$ and $\bm{x}^{+}(\cdot,t)\in L^{2}([0,1])^{n^{+}}$ denote the boundary system’s as well as the backward and forward transport systems’ states. This model is also known as a PDE-ODE system in the backstepping literature, cf. i.e. [8].
For (15), we assume decoupled PDEs with $\bm{\Lambda}=\mathrm{diag}(\mathbf{\Lambda}^{-},-\mathbf{\Lambda}^{+})$ where $\mathbf{\Lambda}^{\pm}=\mathrm{diag}(\lambda_{1}^{\pm},\dotsc,\lambda_{n^{\pm}}^{\pm})$, $\lambda_{i}^{\pm}>\lambda_{i+1}^{\pm}>0$, the pair $(\bm{F},\bm{B})$ controllable and $\operatorname{rank}\bm{Q}_{0}={n^{+}}$ as well as $\operatorname{rank}\bm{Q}_{1}={n^{-}}$.
Figure 1: Structure of the assumed system class
IV-B Parametrisation
As (15a) is controllable, a flatness-based parametrisation in the form of
$\bm{\xi}(t)$ $=\mathbf{N}(\partial_{t})\bm{y}(t)$ (16a)
$\bm{x}^{-}(0,t)$ $=\mathbf{D}(\partial_{t})\bm{y}(t)$ (16b)
can always be given. Furthermore, if the components of the flat output $\bm{y}(t)$ are chosen as the outputs of each block’s integrator chain of the systems’ MIMO controller form (2), the parametrisation matrix $\mathbf{D}(\partial_{t})$ is column reduced and $\operatorname{coldeg}_{i}\mathbf{N}(\partial_{t})<\operatorname{coldeg}_{i}\mathbf{D}(\partial_{t}),\;i\in\{1,\dotsc,m\}$ holds for the column degrees.
Regarding the PDE subsystem, with $\bm{\xi}(t)$ as well as $\bm{x}^{-}(0,t)$ known, (15c) yields $\bm{x}^{+}(0,t)=\left(\mathbf{Q}_{0}\mathbf{D}(\partial_{t})+\mathbf{C}\mathbf{N}(\partial_{t})\right)\bm{y}(t)$ and thus the boundary values required to solve (15b) up to $z=1$. This yields
$\bm{x}^{-}(1,t)$ $=\mathbf{P}(\sigma,t)\bm{x}^{-}(0,t)$ (17)
$\bm{x}^{+}(1,t)$ $=\mathbf{\Delta}(\sigma,\partial_{t})\bm{x}^{+}(0,t)$ (18)
with the prediction matrix $\mathbf{P}(\sigma)=\operatorname{diag}(\sigma^{\tau_{1}^{-}},\dotsc,\sigma^{\tau_{{n^{-}}}^{-}})$ and the delay matrix $\mathbf{\Delta}(\sigma)=\operatorname{diag}(\sigma^{-\tau_{1}^{+}},\dotsc,\sigma^{-\tau_{{n^{+}}}^{+}})$ where $\tau_{i}^{\pm}=\frac{1}{\lambda_{i}^{\pm}}$. Finally, combining Equations 16a, 16b, 17 and 18 with (15d) yields the matrix
$\mathbf{H}(\sigma,\partial_{t})\!=\!\mathbf{P}(\sigma)\mathbf{D}(\partial_{t})-\mathbf{Q}_{1}\mathbf{\Delta}(\sigma)\left(\mathbf{Q}_{0}\mathbf{D}(\partial_{t})+\mathbf{C}\mathbf{N}(\partial_{t})\right)$ (19)
for the parametrisation
$\bm{u}(t)=\mathbf{H}(\sigma,\partial_{t})\bm{y}(t)$ (20)
of the system input by the flat output of the ODE.
As was already to be expected from the motivational example’s parametrisation (10), (19) now structurally shows that $\mathbf{\mathbf{P}}(\sigma)$ and $\mathbf{\Delta}(\sigma)$ introduce all predictions and delays that occur in the system into every column of $\mathbf{H}(\sigma,\partial_{t})$ which poses a problem. Next, we will extend the notion of polynomials in $\partial_{t}$ to incorporate shifts and remedy this inconvenience.
IV-C Generalised polynomials
With the set $T=\{\tau_{1}^{-},\dotsc,\tau_{n^{-}}^{-},\tau_{1}^{+},\dotsc,\tau_{n^{+}}^{+},\}$ of all unique predictions and shifts in the system, by interpreting the members of $\Sigma=\{\sigma^{\tau},\sigma^{-\tau}|\tau\in T\}$ and $\partial_{t}$ as independent generators, we conceive the entries $h_{i,j}(\sigma,\partial_{t})$ of $\mathbf{H}(\sigma,\partial_{t})$ to be from $\mathcal{R}_{\mathrm{p}}=\mathbb{R}[\Sigma,\partial_{t}]$, the space of generalised polynomials in $\sigma$ with real exponents, whose coefficients are polynomials in $\partial_{t}$ with coefficients from $\mathbb{R}$. Furthermore, for the elements
$r(\sigma,\partial_{t})=\sum_{i=1}^{n}a_{i}(\partial_{t})\sigma^{\alpha_{i}}\in\mathcal{R}_{\mathrm{p}}$ (21)
with $a_{i}(\partial_{t})\in\mathbb{R}[\partial_{t}]$, the ordering $\alpha_{i}>\alpha_{i+1}$, for $i\in\{1,\dotsc,n-1\}$ is assumed. To ease working on the polynomials (21), we define the pseudo degrees
$\deg_{\sigma}^{+}r(\sigma,\partial_{t})\coloneq\max_{i\in\{1,\dotsc,n\}}\alpha_{i},\;\deg_{\sigma}^{-}r(\sigma,\partial_{t})\coloneq\min_{i\in\{1,\dotsc,n\}}\alpha_{i}$
as well as the degree [^4] This choice is motivated by the case of Laurent polynomials, which posses positive and negative integer exponents and form an euclidean ring for the given degree definition, cf. [9, 10].
$\deg_{\sigma}r(\sigma,\partial_{t})\coloneq\deg_{\sigma}^{+}r(\sigma,\partial_{t})-\deg_{\sigma}^{-}r(\sigma,\partial_{t}),$ (22)
yielding the upper and lower boundary as well as the diameter of the interval, respectively. Next, we discuss how these intervals can be reduced in a systematic way.
IV-D Shift interval reduction
IV-D1 Preliminary Transformations
Lemma 1
By permuting the input and/or flat output, $\mathbf{H}(\sigma,\partial_{t})$ can be sorted, such that the degrees w.r.t. $\sigma$ in each column are ascending and the highest degrees w.r.t. $\partial_{t}$ appear on the main diagonal, thus with $j\in\{1,\dotsc,m\}$
$\deg_{\sigma}{h}_{i,j}(\sigma,\partial_{t})$ $\leq\deg_{\sigma}{h}_{i+1,j}(\sigma,\partial_{t}),$ (23a)
$\deg_{\partial_{t}}{h}_{l,j\neq}(\sigma,\partial_{t})$ $\leq\deg_{\partial_{t}}{h}_{l,l}(\sigma,\partial_{t})$ (23b)
hold for $i\in\{1,\dotsc,m-1\}$ and $l\in\{1,\dotsc,m\}$.
Proof:
As from (19), there are equal predictions and delays in each row of $\mathbf{H}(\sigma,\partial_{t})$ due to the diagonal structure of $\mathbf{P}(\sigma)$ and $\mathbf{\Delta}(\sigma)$. Then, by (22) the same follows for their corresponding degrees which is why the ordering (23a) is easily obtained by permuting the flat output.
Now, from the parametrisation (16) we have that $\operatorname{coldeg}\mathbf{N}(\partial_{t})<\operatorname{coldeg}\mathbf{D}(\partial_{t})$, which is why $\mathbf{C}$ has no effect regarding the degrees of $\partial_{t}$} in (19) and also that $\mathbf{D}(\partial_{t})$ is column reduced. Since this property is preserved under row transformations over $\mathbb{R}$ and $\mathbf{Q}_{0},\mathbf{Q}_{1}$ have full rank, we obtain that $\mathbf{H}(\sigma,\partial_{t})$ is column reduced w.r.t. $\partial_{t}$. Finally, as the elements in each row have equal degree w.r.t. $\sigma$, (23b) is achieved by permuting the input. ∎
IV-D2 Goal
As discussed in the motivation, to obtain the controller form (12), the input parametrisation (20) must be solved for the highest derivatives of the flat output. Furthermore, to introduce a HCF state in a direct fashion, these must occur together with the largest predictions as well as largest delays. In the algebraic setting, this corresponds to the composition of taking the LCCMs w.r.t. $\sigma$ and $\partial_{t}$ of $\mathbf{H}(\sigma,\partial_{t})$ to commute:
$\operatorname{lccm}_{s}(\operatorname{lccm}^{+}_{\sigma}\mathbf{H}(\sigma,\partial_{t}))$ $=\operatorname{lccm}^{+}_{\sigma}(\operatorname{lccm}_{s}\mathbf{H}(\sigma,\partial_{t}))$ (24a)
$\operatorname{lccm}_{s}(\operatorname{lccm}_{\sigma}\mathbf{H}(\sigma,\partial_{t}))$ $=\operatorname{lccm}_{\sigma}(\operatorname{lccm}_{s}\mathbf{H}(\sigma,\partial_{t}))$ (24b)
and the result having full rank. The condition for $\deg_{\sigma}^{-}$ then follows from (22). These requirements are trivially satisfied if the degree conditions
$\deg^{+}_{\sigma}{h}_{i,j\neq i}(\sigma,\partial_{t})$ $\leq\deg^{+}_{\sigma}{h}_{i,i}(\sigma,\partial_{t})$ (25a)
$\deg_{\sigma}{h}_{i,j\neq i}(\sigma,\partial_{t})$ $\leq\deg_{\sigma}{h}_{i,i}(\sigma,\partial_{t})$ (25b)
$\deg_{\partial_{t}}{h}_{i,j\neq i}(\sigma,\partial_{t})$ $\leq\deg_{\partial_{t}}{h}_{i,i}(\sigma,\partial_{t})$ (25c)
hold for all $i,j\in\{1,\dotsc,{m}\}$, since all LCCMs in this case are diagonal. Furthermore, as by Lemma 1, (25a),(25b) for $j\leq i$ as well as (25c) for $j\in\{1,\dotsc,m\}$ are satisfied, the elements $h_{i,i}(\sigma,\partial_{t})$ of each column are now used to reduce the (pseudo) degrees of the entries below them, such that all
$\bar{h}_{j,i}(\sigma,\partial_{t})=h_{i,j}(\sigma,\partial_{t})-q_{j,i}(\sigma,\partial_{t})h_{i,i}(\sigma,\partial_{t})$ (26)
for $i,j\in\{1,\dotsc,m\},j>i$ satisfy (25). Next, we discuss a generalised polynomial long division (GPLD) algorithm to perform such a reduction for a single element.
IV-D3 Generalised polynomial long division
For now, we treat all entries in (26) as elements of $\mathcal{R}_{\mathrm{r}}=\mathbb{R}(\partial_{t})\left[\Sigma\right]$ which allows for rational expressions in $\partial_{t}$ as coefficients and therefore enables us to develop a polynomial long division scheme regarding $\sigma$.
Lemma 2
Let $x(\sigma,\partial_{t}),y(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{r}},\;y(\sigma,\partial_{t})\neq 0$ be generalised polynomials with rational coefficients where $\deg_{\sigma}x(\sigma)\geq\deg_{\sigma}y(\sigma)$. Then there exist $q(\sigma,\partial_{t}),r(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{r}}$ such that
$x(\sigma,\partial_{t})=q(\sigma,\partial_{t})y(\sigma,\partial_{t})+r(\sigma,s)$ (27)
with $\deg_{\sigma}r(\sigma,\partial_{t})<\deg_{\sigma}y(\sigma,\partial_{t})$ as well as $\deg^{+}_{\sigma}r(\sigma,\partial_{t})<\deg^{+}_{\sigma}y(\sigma,\partial_{t})$ and $\deg^{-}_{\sigma}r(\sigma,\partial_{t})>\deg^{-}_{\sigma}y(\sigma,\partial_{t})$.
Proof:
As the following steps are valid for arbitrary coefficients, the explicit mention of $\partial_{t}$ will be omitted. In the general case we have $\deg_{\sigma}x(\sigma)>\deg_{\sigma}y(\sigma)$ as well as $\deg^{+}_{\sigma}x(\sigma)>\deg^{+}_{\sigma}y(\sigma)$, $\deg^{-}_{\sigma}x(\sigma)<\deg^{-}_{\sigma}y(\sigma)$ or both. Since the exponents of the polynomials $x(\sigma,\partial_{t})=\sum_{i=1}^{n}a_{i}(\partial_{t})\sigma^{\alpha_{i}}$, $y(\sigma,\partial_{t})=\sum_{i=1}^{m}b_{i}(\partial_{t})\sigma^{\beta_{i}}$ are ordered, we use
$q(\sigma)=\begin{cases}\frac{a_{1}\sigma^{\alpha_{1}}}{b_{1}\sigma^{\beta_{1}}},&\text{ if }\deg_{\sigma}^{+}x(\sigma)>\deg^{+}_{\sigma}y(\sigma)\\ \frac{a_{n}\sigma^{\alpha_{n}}}{b_{m}\sigma^{\beta_{m}}},&\text{ if }\deg^{-}_{\sigma}x(\sigma)<\deg^{-}_{\sigma}y(\sigma),\end{cases}$ (28)
where in the case that both are true, the upper branch is taken. Assuming the upper branch was taken, this yields
$\begin{split}r(\sigma)=x(\sigma)-q(\sigma)y(\sigma)=\sum\limits_{i=2}^{n}a_{i}\sigma^{\alpha_{i}}-\frac{a_{1}\sigma^{\alpha_{1}}}{b_{1}\sigma^{\beta_{1}}}\sum\limits_{i=2}^{m}b_{i}\sigma^{\beta_{i}},\end{split}$
for which we find that
$\deg^{+}_{\sigma}r(\sigma)=\max\left\{\alpha_{2},\alpha_{1}-(\beta_{1}-\beta_{2})\right\}<\deg^{+}_{\sigma}x(\sigma)$
as well as $\deg^{-}_{\sigma}r(\sigma)=\min\{\alpha_{n},\alpha_{1}-(\beta_{1}-\beta_{m})\}$ in which we identify $\deg_{\sigma}y(\sigma)=\beta_{1}-\beta_{m}$ and use $\alpha_{1}=\deg_{\sigma}x(\sigma)+\alpha_{n}$ as well as $\kappa=\deg_{\sigma}x(\sigma)-\deg_{\sigma}y(\sigma)\geq 0$ to obtain
$\begin{split}\deg^{-}_{\sigma}r(\sigma)=\min\{&\alpha_{n},\alpha_{n}+\kappa\}=\alpha_{n}=\deg^{-}_{\sigma}x(\sigma).\end{split}$
This shows that the predictions in $x(\sigma)$ can be reduced without affecting the delays. Regarding the overall degree, we have $\deg_{\sigma}r(\sigma)=\max\{\alpha_{2}-\alpha_{n},\alpha_{1}-\alpha_{n}-(\beta_{1}-\beta_{2})\}$ which yields $\deg_{\sigma}r(\sigma)<\deg_{\sigma}x(\sigma)$. Hence, $x(\sigma)$ can be reduced in a manner such that the iteration
$r_{k}(\sigma)=r_{k-1}(\sigma)-q_{k}(\sigma)y(\sigma)$
with $r_{0}(\sigma)=x(\sigma)$ and $q_{k}(\sigma)=\frac{c_{k,1}\sigma^{\gamma_{k,1}}}{b_{1}\sigma^{\beta_{1}}}$ by $r_{k}(\sigma)=\sum_{i=1}^{n}c_{k,i}\sigma^{\gamma_{k,i}}$ will for some $k^{+}$ terminate with $\deg^{+}_{\sigma}r_{k^{+}}(\sigma)<\deg^{+}_{\sigma}y(\sigma)$ since $x(\sigma)$ only has a finite amount of summands.
Now, (if required) the second branch of (28) is taken and similar computations yield that restarting the iteration with $r_{0}(\sigma)=r_{k^{+}}(\sigma)$ and choosing $q_{k}(\sigma)=\frac{c_{k,1}\sigma^{\gamma_{k,1}}}{b_{m}\sigma^{\beta_{m}}}$ will yield $\deg^{-}_{\sigma}r_{k^{-}}(\sigma)>\deg^{-}_{\sigma}y(\sigma)$ after $k^{-}$ more steps. Summing up all $q_{k}(\sigma)$ from both iterations then yields the resulting “quotient” $q(\sigma)$ with common denominator $d=b_{1}^{k^{+}}b_{m}^{k^{-}}$. ∎
With Lemma 2, the shift degrees are taken care of. However, the coefficients of $q(\sigma,\partial_{t}),r(\sigma,\partial_{t})$ are rational expressions in $\partial_{t}$. Furthermore, in the analytic setting, their denominators can be interpreted as functions that their numerator polynomials are applied to. This results in piecewise continuous functions which do not necessarily have compact support. In other words: They generally do not correspond to differential or (distributed) delay operators with finite shift amplitude from $\mathcal{E}^{\prime}$. However, thanks to the Palay-Wiener theorem, if their Laplace transforms are entire functions, they indeed are such operators, cf. [11, Thm. 7.3.1].
Lemma 3
Let $x(\sigma,\partial_{t}),y(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{p}},\;y(\sigma,\partial_{t})\neq 0$ be generalised polynomials with polynomial coefficients where $\deg_{\sigma}x(\sigma)\geq\deg_{\sigma}y(\sigma)$, as well as $\deg_{\partial_{t}}x(\sigma)\leq\deg_{\partial_{t}}y(\sigma)$. Then there exist $q^{\star}(\sigma,\partial_{t}),r^{\star}(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{w}}=\mathcal{R}_{\mathrm{r}}\cap\mathcal{E}^{\prime}$ such that
$x(\sigma,\partial_{t})=q^{\star}(\sigma,\partial_{t})y(\sigma,\partial_{t})+r^{\star}(\sigma,s)$ (29)
with $\deg_{\sigma}r^{\star}(\sigma,\partial_{t})=\deg_{\sigma}y(\sigma,\partial_{t})$, $\deg^{+}_{\sigma}r^{\star}(\sigma,\partial_{t})=\deg^{+}_{\sigma}y(\sigma,\partial_{t})$ and $\deg^{-}_{\sigma}r^{\star}(\sigma,\partial_{t})=\deg^{-}_{\sigma}y(\sigma,\partial_{t})$.
Proof:
This proof is heavily inspired by [12, 9] and only adapted to the current setup. First, we obtain $q(\sigma,\partial_{t}),r(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{r}}$ by means of Lemma 2 and extract its normalised common denominator $d(\partial_{t})\in\mathbb{R}[\partial_{t}]$ such that $q(\sigma,\partial_{t})=\frac{\tilde{q}(\sigma,\partial_{t})}{d(\partial_{t})}$ with $\tilde{q}(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{p}}$. Next, the factorisation $d(\partial_{t})=\prod_{k=1}^{D}(\partial_{t}-s_{k})$ with the roots $s_{k}$ and $D=\deg d(\partial_{t})$ are used together with $\hat{q}_{0}(s)=\mathcal{L}\{\tilde{q}(\sigma,\partial_{t})\}$ in the iteration
$\hat{q}_{k}(s)=\frac{\hat{q}_{k-1}(s)-\hat{q}_{k-1}(s_{k})}{s-s_{k}},$ (30)
which introduces a nominator zero for every pole from $q(\sigma,\partial_{t})$ it adds. Thus, $q_{D}(s)$ obtained after $D$ iterations is an entire function on $\mathbb{C}$ and therefore $q^{\star}(\sigma,\partial_{t})=\mathcal{L}^{-1}\{q_{D}(s)\}\in\mathcal{R}_{\mathrm{w}}$ by [11, Thm. 7.3.1].
Now, due to the structure of (30) we can justify the separation $q^{\star}(\sigma,\partial_{t})=q(\sigma,\partial_{t})+p(\partial_{t})$ with $p(\partial_{t})\in\mathbb{R}(\partial_{t})$. Therefore, we obtain the degree conditions from Lemma 2 with equal signs for the new remainder $r^{\star}(\sigma,\partial_{t})$. ∎
IV-D4 Algorithm
Equipped with a suitable way to reduce the degrees of its elements by means of the GPLD from Lemma 3, we now reduce the degrees in the entire matrix $\mathbf{H}(\sigma,\partial_{t})$ by applying a series of row transformations.
Algorithm 1 Shift reduction in $\mathbf{H}(\sigma,\partial_{t})$
alg]alg:shift_red
1: $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]function compute_transform($\mathbf{X},j$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
2:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\mathbf{L}_{c}\leftarrow\mathbf{I}_{{n^{-}}}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
3:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]for $i\in\{j+1,\dotsc,{n^{-}}\}$ do $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
4:   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] // From Lemma 3: $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
5:   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $q^{\star}_{i,j}\leftarrow$ GPLD($x=\mathbf{X}[i,j],y=\mathbf{X}[j,j]$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
6:   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\mathbf{L}_{c}[i,j]\leftarrow-q^{\star}_{i,j}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
7:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] return $\mathbf{L}_{c}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
8:
9: $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]procedure reduce_shifts($\mathbf{H}$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
10:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\bar{}\mathbf{H}\leftarrow\mathbf{H}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
11:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]for $i\in\{2,\dotsc,{n^{-}}\}$ do $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
12:   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]for $j\in\{1,\dotsc,{n^{-}}-1\}$ do $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
13:     $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\mathbf{L}_{c}\leftarrow$ Compute transform($\bar{}\mathbf{H},j$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
14:     $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\bar{}\mathbf{H}\leftarrow\mathbf{L}_{c}\mathbf{H}$, $\mathbf{L}\leftarrow\mathbf{L}_{c}\mathbf{L}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
15:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] return $\bar{}\mathbf{H}$$\mathbf{L}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
Lemma 4
The matrix $\bar{}\mathbf{H}(\sigma,\partial_{t})=\mathbf{L}(\sigma,\partial_{t})\mathbf{H}(\sigma,\partial_{t})$ obtained by LABEL:alg:shift_red satisfies the requirements (24).
(25).
Proof:
Starting in the second row, the degrees in each column are corrected from left to right, where each of these correction steps will introduce new undesired shifts to its left in the rows below its diagonal element. Therefore, the procedure then continues to the next row, until no more new shifts are introduced as the last row is reached. ∎
IV-E Hyperbolic MIMO controller form
Applying LABEL:alg:shift_red to $\mathbf{H}(\sigma,\partial_{t})$, we obtain
$\mathbf{L}^{-1}(\sigma,\partial_{t})\bm{u}(\partial_{t})=\bar{}\mathbf{H}(\sigma,\partial_{t})\bm{y}(\partial_{t})$ (31)
with $\hat{\mathbf{H}}=\operatorname{lccm}_{s,\sigma}\bar{}\mathbf{H}(\sigma,\partial_{t})$ having full rank. Thus, by means of the separation $\bar{}\mathbf{H}(\sigma,\partial_{t})=\hat{}\mathbf{H}\mathbf{K}(\sigma,\partial_{t})+\tilde{}\mathbf{H}(\sigma,\partial_{t})$ where $\mathbf{K}(\sigma,\partial_{t})=\operatorname{diag}(\partial_{t}^{\nu_{1}}\sigma^{\hat{\tau}_{1}},\dotsc,\partial_{t}^{\nu_{m}}\sigma^{\hat{\tau}_{m}})$ with the highest derivative and prediction orders $\nu_{i}=\deg_{\partial_{t}}\bar{h}_{i,i}(\sigma,\partial_{t})$ and $\hat{\tau}_{i}=\deg^{+}_{\sigma}\bar{h}_{i,i}(\sigma,\partial_{t})$ we obtain the explicit DDE system
$\mathbf{K}(\sigma,\partial_{t})\bm{y}(\partial_{t})=\hat{\mathbf{H}}^{-1}\!\left(\mathbf{L}^{-1}(\sigma,\partial_{t})\bm{u}(\partial_{t})\!-\tilde{\mathbf{H}}(\sigma,\partial_{t})\bm{y}(\partial_{t})\right)$ (32)
which reads
$y_{i}^{(\nu_{i})}(t+\hat{\tau}_{i})=-\mathcal{A}_{i}(t)*\bm{y}(t)+\mathcal{B}_{i}(t)*\bm{u}(t)$ (33)
for $i=\{1,\dotsc,m\}$ in the analytic setting. Herein, the elements of $\mathcal{A}_{i},\mathcal{B}_{i}$ by construction have compact support, which is why the convolutions will yield definite integrals whose intervals depend on the degrees of $\tilde{\mathbf{H}}(\sigma,\partial_{t})$ and $\mathbf{L}^{-1}(\sigma,\partial_{t})$ w.r.t $\sigma$. Thus, (33) states a well posed initial value problem if the components $y_{i}(t+\tau_{i})$ of the flat output are known for $\tau_{i}\in[\check{\tau}_{i},\hat{\tau}_{i}]$ with $\check{\tau}_{i}=\deg^{-}_{\sigma}\bar{h}_{i,i}(\sigma,\partial_{t})$ and $\bm{u}$ is known on $\mathbb{R}$.
Hence, we introduce the HCF state
$\eta^{\mathrm{O}}_{i,j}(t)$ $\coloneq y_{i}^{(j-1)}(t-\check{\tau}_{i}),\;j\in\{1,\dotsc,\nu_{i}\}$ (34a)
$\eta^{\mathrm{P}}_{i}(\tau,t)$ $\coloneq y_{i}^{(\nu_{i})}(t+\tau),\;\tau\in[\check{\tau}_{i},\hat{\tau}_{i}]$ (34b)
for $i\in\{1,\dotsc,m\}$ which is equivalent to the flat output on these intervals. Substituting [^5] Some expressions in $\bm{y}(t)$ may remain as e.g. the values $y_{i}^{(d)}(t+\tau_{i})$ are not part of the HCF state $\bm{\eta}(\cdot,t)$ for $p<\nu_{i}$, cf. (12). However, these expressions can be easily eliminated by means shown e.g. in [6, Lemma 5]. the definitions (34) in (33) then allows to identify the functions $a^{\star}_{i,j}(t),b^{\star}_{i,j}(t)$ as well as weights $\bm{a}_{i}$ in (12), yielding the HCF of system (15).
Lemma 5
The type of HCF obtained is a quasi controller form, (cf. Section III-C3), with only input predictions and delays but no input derivatives.
Proof:
The proof of Lemma 2 yields $\deg_{\partial_{t}}q(\sigma,\partial_{t})<0$ due to the structure of $d(\partial_{t})$. Then by Lemma 3, $q^{\star}(\sigma,\partial_{t})$ has non-positive degree as do all entries in $\mathbf{L}_{i}(\sigma,\partial_{t})$ from LABEL:alg:shift_red. Now, as the inverse of the triangular $\mathbf{L}(\sigma,\partial_{t})$ is computed by forward substitution, all $\bar{q}_{i,j}(\sigma,\partial_{t})$ therein are linear combinations of the $q_{i,j}^{\star}(\sigma,\partial_{t})$, hence, $\deg_{\partial_{t}}\bar{q}_{i,j}(\sigma,\partial_{t})\leq 0$. Therefore, the convolution with $\mathcal{B}_{i}$ yields no derivatives. ∎
V Reduction of the benchmark system
Consider again the system from Section II-C with $\alpha=\pi$, $\beta=10$ and $m=1$ substituted into (10) yielding
$\mathbf{H}(\sigma,\partial_{t})=\frac{1}{4}\begin{pmatrix}a_{1}(\partial_{t})\sigma^{\pi}+a_{2}(\partial_{t})\sigma^{-\pi}&-\sigma^{\pi}-\sigma^{-\pi}\\ b_{1}(\partial_{t})\sigma^{10}+b_{2}(\partial_{t})\sigma^{-10}&\sigma^{10}+\sigma^{-10}\\ \end{pmatrix}$
with $a_{1}(\partial_{t})=b_{1}(\partial_{t})=\partial_{t}^{2}+2\partial_{t}$ and $a_{2}(\partial_{t})=b_{2}(\partial_{t})=\partial_{t}^{2}-2\partial_{t}$ for which $\operatorname{lccm}_{\sigma,\partial_{t}}\mathbf{H}(\sigma,\partial_{t})=\operatorname{lccm}_{\partial_{t},\sigma}\mathbf{H}(\sigma,\partial_{t})$ is of rank $1$ where $2$ would be needed.
Now, regarding Section IV-D1, the rows are already correctly ordered since $\pi<10$ and also the columns do not need to be reordered since the coefficients in each column are of the same degree w.r.t. $\partial_{t}$. Continuing with Section IV-D2, the conditions (25) fail for $h_{2,1}(\sigma,\partial_{t})$ as e.g. $\deg_{\sigma}h_{2,1}(\sigma,\partial_{t})=20>\deg_{\sigma}h_{1,1}(\sigma,\partial_{t})=2\pi$. Therefore, by means of Lemma 2 the quotient
$\begin{aligned} q_{1,2}&(\sigma,\partial_{t})\!=\!\sigma^{10-\pi}\!+\!\frac{2-\partial_{t}}{\partial_{t}+2}\sigma^{10-3\pi}\!-\!\frac{\partial_{t}+2}{\partial_{t}-2}\sigma^{-10+3\pi}\!+\!\sigma^{-10+\pi}\end{aligned}$
is obtained, where one iteration was needed for each case in (28). Next, Lemma 3 yields the correction term $p(\partial_{t})=\frac{16e^{-20+6\pi}}{\left(\partial_{t}^{2}-4\right)}$ and thus $q_{2,1}^{\star}(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{w}}$. Using the latter in (26) then yields the reduced element
$\begin{aligned} \bar{h}_{1,2}(\sigma,\partial_{t})=\!-\frac{16\partial_{t}}{\left(\partial_{t}-2\right)e^{20-6\pi}}\sigma^{\pi}\!+\frac{\partial_{t}\left(\partial_{t}^{2}+4\partial_{t}+4\right)}{\partial_{t}-2}\sigma^{-10+4\pi}\\ \!+\frac{\partial_{t}\left(\partial_{t}^{2}-4\partial_{t}+4\right)}{\partial_{t}+2}\sigma^{10-4\pi}\!-\frac{16\partial_{t}}{\left(\partial_{t}+2\right)e^{20-6\pi}}\sigma^{-\pi}.\end{aligned}$
As the requirements for the second column are already satisfied, this concludes the reduction and we obtain $\hat{}\mathbf{H}=\operatorname{lccm}_{\partial_{t},\sigma}\bar{}\mathbf{H}(\sigma,\partial_{t})=\operatorname{diag}(1,-2)$.
Python code showing the complete computation of the HCF will be made available at a later time.
VI Concluding remarks
This contribution establishes a HCF for MIMO systems with dynamic boundaries and proposes a flatness based algorithm to compute said form for a specific class of systems. While the algebraic approach taken to obtain this result is limited to linear systems, it enables a systematic procedure which lends well to an implementation using computer algebra. However, the insight gained from an analytic interpretation of the algebraic procedure allows for the technique to be extended to nonlinear boundary systems, which will be discussed in a forthcoming publication. Furthermore, the obtained HCF contains input delays, which will result in a dynamic extension, thus further analysis is required to obtain a form without delays as this would enable the application of quasi-static state feedback. Finally, the system class (15) is lacking in-domain couplings, as these were recently addressed in [13] for the MIMO case without boundary dynamics, the next logical step is to combine the two approaches.

### v3

@@
 This contribution develops an algebraic approach to obtain a controller form for a class of linear hyperbolic MIMO systems, bidirectionally coupled with a linear ODE system at the unactuated boundary.
-After a short summary of established controller forms for SISO and MIMO ODE as well as SISO hyperbolic PDE systems, it is shown that the direct approach to state a controller form fails already for a very simple MIMO example.
-Next, a generalised hyperbolic controller form with different variants is proposed and a new flatnesss-based scheme to compute said form is presented.
-Therein, the system is treated in an algebraic setting where generalised polynomials with real exponents are used to describe the predictions and delays in the system.
+After a short summary of established controller forms for SISO and MIMO ODE as well as SISO hyperbolic PDE systems, it is shown that the approach to state a controller form for SISO systems cannot easily be transferred to the MIMO case as it already fails for a very simple example.
+Next, a generalised hyperbolic controller form with different variants is proposed and a new flatness-based scheme to compute said form is presented.
+Therein, the system is treated in an algebraic setting where quasipolynomials are used to express the predictions and delays in the system.
 The proposed algorithm is then applied to the motivating example.
-©2025 IEEE.
+©2026 IEEE.
 Personal use of this material is permitted.
@@
 I Introduction
-From the onset of state space based approaches, controller (canonical) forms have been a vital asset for systematic controller designs.
-As the controller forms for linear SISO and MIMO ODE systems, cf.
-[1, 2] as well for SISO hyperbolic PDE systems [3] are long established and understood, only comparatively recent works [4, 5, 6] discuss hyperbolic systems with boundaries governed by ODEs for the SISO case which this contribution extends to the MIMO case.
-Following some preliminaries, where we briefly restate established controller forms and demonstrate the problems that arise if a hyperbolic controller form is to be introduced even for the most simple MIMO system, a MIMO hyperbolic controller form (HCF) is given in Section III.
-The main part of the contribution in Section IV then discusses how a HCF can be obtained for a specific class of systems whereupon Section V tracks the procedure for an example system.
+Canonical forms serve as a cornerstone for systematic feedback design in finite dimensions.
+Their counterparts for distributed parameter systems (DPSs) have been studied since the late 1970s.
+For linear hyperbolic partial differential equations (PDEs), [1] connected controllability and spectral assignment to controller canonical forms (CCFs) and neutral delay equations.
+More recent works linked the aforementioned delay interpretations to flatness based parametrisations [2, 3] (as well as [4] for early ideas in this direction) and established the connection between such parametrisations and the CCF for more general systems.
+In particular, [5] introduced a hyperbolic controller canonical form (HCCF) for linear hyperbolic SISO systems comprising general interconnections of hyperbolic PDEs and ordinary differential equations (ODEs) (see also [6]) by generalising the familiar chain-of-integrators structure by prepending a transport equation to represent the delays of the system imposed by the hyperbolic PDEs.
+Therein, the canonical coordinates are deduced from the feasible initial conditions (ICs) for a functional differential equation (FDE) arising from a flatness-based parametrisation of the system’s input while the required state transforms are computed using the flatness-based parametrisation of the system’s original state variables.
+Aside from the above-mentioned state-space approaches, in [7, 8] flatness-based parametrisations also served as means for purely algebraic analysis and control design.
+These methods employ and partially generalise results from [9] and [10].
+As many relevant technological applications have multiple in- and outputs, several control approaches are available, e.g.
+[11] using flatness-based methods or [12] employing backstepping designs.
+However, regarding controller forms a direct extension of the existing work to the MIMO case is difficult, as such systems can exhibit different delays depending on the considered input which together with the bidirectionally coupled ODE result in an FDE system for which feasible ICs are not obvious, thus, negating the approach from the SISO case.
+Therefore, the goal of this contribution is to introduce a hyperbolic controller form (HCF) for a class of linear hyperbolic MIMO systems with dynamic boundary conditions and to develop a delay reduction algorithm such that the resulting FDE allows for a direct computation of the controller form.
+To do so, we denote the input-output relation using quasipolynomials, whose use in the control of hyperbolic systems was already discussed in [13, 14], and restate the requirements to introduce a HCF as column-reduced properties on the resulting quasipolynomial matrices.
+Next, in contrast to the established methods and due to the incommensurate delays in the considered systems as well as the interplay with the boundary system, we propose a modified quasipolynomial division algorithm to compute the required transformation to establish said properties.
+Following some preliminaries, where we briefly restate established controller forms and demonstrate the problems that arise if a hyperbolic controller form is to be introduced even for the most simple MIMO system, a general MIMO HCF and several subforms are given in Section III.
+The main part of the contribution in Section IV then discusses how a HCF can be obtained for a specific class of systems Finally, Section V traces the procedure for an example system.
 II Preliminaries
@@
 The first and second derivative of $x(t)$ is denoted by $\dot{x}(t)$ and $\ddot{x}(t)$, while the $n$-th derivative is given by $x^{(n)}(t)$ and the partial derivative of $x(z,t)$ w.r.t.
-$z$.
-by $\partial_{z}x(z,t)$.
-To treat delay differential equations (DDEs) in an algebraic setting, we identify the operators therein with their Fourier-Laplace transforms and treat the application of the transformed operators as convolutions.
-Thus, we write $\mathcal{L}\{\partial_{t}\}=s$ for differentials, while the shifts $(\sigma^{\tau}x)(t)=x(t+\tau)$ with $\tau\in\mathbb{R}$ are denoted as $\mathcal{L}\{\sigma^{\tau}\}(s)=\exp(s\tau)$.
+$z$ by $\partial_{z}x(z,t)$.
+To treat FDEs in an algebraic setting, we identify the operators therein with their Fourier-Laplace transforms and treat the application of the transformed operators as convolutions.
+Thus, we write $\mathcal{L}\{{\mathrm{d}_{t}}\}=s$ for differentials, while the shifts $(\sigma^{\tau}x)(t)=x(t+\tau)$ with $\tau\in\mathbb{R}$ are denoted as $\mathcal{L}\{\sigma^{\tau}\}(s)=\exp(s\tau)$.
 The shorthand $L^{2}([a,b])$ is used for the space of square integrable functions on $[a,b]$, while $H^{n}([a,b])$ stands for the space of functions whose $n$-th derivative belongs to $L^{2}([a,b])$ and $\mathcal{E}^{\prime}$ for Schwartz-Distributions with compact support.
-With $\operatorname{V}_{f}([a,b])$ denoting the total variation of a function $f$ on the interval $[a,b]$, $BV([a,b])$ contains functions of bounded variation while members of its subset $BV^{\star}([a,b])$ satisfy $\lim_{\epsilon\to 0}\operatorname{V}_{f}([b,b+\epsilon))=0$.
-Furthermore, $\mathbb{R}[\partial_{t}]$ is the space of polynomials in $\partial_{t}$ with coefficients in $\mathbb{R}$ and $\mathbb{R}(\partial_{t})$ consists of rational expressions in $\partial_{t}$.
-Regarding a polynomial matrix $X(\partial_{t})\in\mathbb{R}^{n\times n}[\partial_{t}]$, we denote the leading column coefficient matrix (LCCM) w.r.t.
-$\partial_{t}$ by $\operatorname{lccm}_{\partial_{t}}{\mathbf{X}(\partial_{t})}$ and say that $\mathbf{X}(\partial_{t})$ is column reduced w.r.t.
-$\partial_{t}$ if $\operatorname{lccm}_{\partial_{t}}{\mathbf{X}(\partial_{t})}$ has full generic rank.
-Finally, we call $\mathbf{X}(\partial_{t})$ unimodular if its determinant is independent of $\partial_{t}$.
+With $V_{f}([a,b])$ denoting the total variation of a function $f$ on the interval $[a,b]$, $\mathrm{BV}([a,b])$ contains functions of bounded variation while members of its subset $\mathrm{BV}^{\star}([a,b])$ satisfy $\lim_{\epsilon\to 0}V_{f}([b,b-\epsilon))=0$.
+Furthermore, $\mathbb{R}[{\mathrm{d}_{t}}]$ is the space of polynomials in $\mathrm{d}_{t}$} with coefficients in $\mathbb{R}$ while $\mathbb{R}({\mathrm{d}_{t}})$ consists of rational expressions in $\mathrm{d}_{t}$}.
+Regarding a polynomial matrix $\mathbf{X}({\mathrm{d}_{t}})\in\mathbb{R}^{n\times n}[{\mathrm{d}_{t}}]$, we denote the leading column coefficient matrix (LCCM) w.r.t.
+$\mathrm{d}_{t}$} by $\operatorname{lccm}_{{\mathrm{d}_{t}}}{\mathbf{X}({\mathrm{d}_{t}})}$ and say that $\mathbf{X}({\mathrm{d}_{t}})$ is column reduced w.r.t.
+$\mathrm{d}_{t}$} if $\operatorname{lccm}_{{\mathrm{d}_{t}}}{\mathbf{X}({\mathrm{d}_{t}})}$ has full generic rank.
+Finally, we call $\mathbf{X}({\mathrm{d}_{t}})$ unimodular over $\mathbb{R}[{\mathrm{d}_{t}}]$ if its determinant is a nonzero constant (and therefore independent of $\mathrm{d}_{t}$}).
 II-B Established controller forms
@@
 $\dot{\eta}^{\mathrm{O}}_{n}(t)$ $=-\bm{a}^{\intercal}\bm{\eta}^{\mathrm{O}}(t)+u(t)$ (1c)
-with state $\bm{\eta}^{\mathrm{O}}(t)=(\eta^{\mathrm{O}}_{1}(t),\dotsc,\eta^{\mathrm{O}}_{n}(t))^{\intercal}\in\mathbb{R}^{n}$, coefficients $\bm{a}\in\mathbb{R}^{n}$ and input $u(t)\in\mathbb{R}$ that exists for every controllable system.
+with state $\bm{\eta}^{\mathrm{O}}(t)=(\eta^{\mathrm{O}}_{1}(t),\dotsc,\eta^{\mathrm{O}}_{n}(t))^{\intercal}\in\mathbb{R}^{n}$, coefficients $\bm{a}\in\mathbb{R}^{n}$ and input $u(t)\in\mathbb{R}$, which exists for every controllable system.
 Note that the “output” $\eta^{\mathrm{O}}_{1}(t)$ of the integrator chain is a flat output of the system.
@@
 $\dot{\eta}^{\mathrm{O}}_{i,\nu_{i}-1}(t)$ $=\eta^{\mathrm{O}}_{i,\nu_{i}}(t)$ (2b)
-$\dot{\eta}^{\mathrm{O}}_{i,\nu_{i}}(t)$ $=-\sum_{k=1}^{m}\bm{a}_{k}^{\intercal}\bm{\eta}^{\mathrm{O}}_{k}(t)+\bm{b}_{i}^{\intercal}\bm{u}(t)$ (2c)
-for $i\in\{1,\dotsc,{m}\}$, with the Kronecker index $\nu_{i}$ and the state vector $\bm{\eta}^{\mathrm{O}}_{i}(t)=(\eta^{\mathrm{O}}_{i,1}(t),\dotsc,\eta^{\mathrm{O}}_{i,\nu_{i}}(t))^{\intercal}\in\mathbb{R}^{\nu_{i}}$ of the $i$-th subsystem.
-Similarly to the SISO case, the MIMO controller form state $\bm{\eta}^{\mathrm{O}}(t)\in\mathbb{R}^{\nu_{\Sigma}}$ with $\nu_{\Sigma}\coloneq\sum_{i=1}^{m}\nu_{i}$ and input $\bm{u}(t)$ can be easily parametrised by the tuple $(\eta^{\mathrm{O}}_{1,1}(t),\dotsc,\eta^{\mathrm{O}}_{m,1}(t))$.
-II-B3 hyperbolic SISO systems
+$\dot{\eta}^{\mathrm{O}}_{i,\nu_{i}}(t)$ $=-\sum_{j=1}^{m}\bm{a}_{ij}^{\intercal}\bm{\eta}^{\mathrm{O}}_{j}(t)+\bm{b}_{i}^{\intercal}\bm{u}(t)$ (2c)
+for $i\in\{1,\dotsc,{m}\}$, with the Kronecker indices $\nu_{i}$ and the state vector $\bm{\eta}^{\mathrm{O}}_{i}(t)=(\eta^{\mathrm{O}}_{i,1}(t),\dotsc,\eta^{\mathrm{O}}_{i,\nu_{i}}(t))^{\intercal}\in\mathbb{R}^{\nu_{i}}$ of the $i$-th subsystem.
+Similarly to the SISO case, the MIMO controller form state $\bm{\eta}^{\mathrm{O}}(t)\in\mathbb{R}^{\nu_{\Sigma}}$ with $\nu_{\Sigma}\coloneq\sum_{i=1}^{m}\nu_{i}$ and input $\bm{u}(t)$ can be easily parametrised by a flat output $(\eta^{\mathrm{O}}_{1,1}(t),\dotsc,\eta^{\mathrm{O}}_{m,1}(t))$.
+Note that, in principle, there exist several ways to distribute the $\nu_{\Sigma}$ integrators over the m{m} chains but only one that yields the form (2) without input derivatives.
+II-B3 Hyperbolic SISO systems with boundary dynamics
 Initially, the work [1] introduced a controller form for boundary controlled hyperbolic systems that takes the form of decoupled transport systems.
@@
 $\dot{\eta}^{\mathrm{O}}_{n-1}(t)$ $=\eta^{\mathrm{O}}_{n}(t)$ (3b)
-$\dot{\eta}^{\mathrm{O}}_{n}(t)$ $=\eta^{\mathrm{P}}(\check{\tau},t)$ (3c)
-[…]
-$\eta^{\mathrm{P}}(\hat{\tau},t)$ $=-\operatorname{\mathcal{S}}_{a^{\star}}\eta^{\mathrm{P}}(\cdot,t)-\bm{a}^{\intercal}\eta^{\mathrm{O}}(t)+u(t).$ (3e)
-[…]
+$\dot{\eta}^{\mathrm{O}}_{n}(t)$ $=\eta^{\mathrm{P}}(0,t)$ (3c)
+$\partial_{t}\eta^{\mathrm{P}}(\tau,t)$ $=\partial_{\tau}\eta^{\mathrm{P}}(\tau,t),\quad\tau\in[0,\hat{}\tau)$ (3d)
+$\eta^{\mathrm{P}}(\hat{}\tau,t)$ $=-\left(a^{\star}\eta^{\mathrm{P}}\right)(t)-\bm{a}^{\intercal}\eta^{\mathrm{O}}(t)+u(t)$ (3e)
+with state $\bm{\eta}(\cdot,t)=(\bm{\eta}^{\mathrm{O}}(t),\eta^{\mathrm{P}}(\cdot,t))\in\mathbb{R}^{n}\times L^{2}([0,\hat{}\tau])$ and the Stieltjes integral $\left(a^{\star}\eta^{\mathrm{P}}\right)(t)=\int_{0}^{\hat{}\tau}\eta^{\mathrm{P}}(\tau,t)\,\mathrm{d}\bar{a}(\tau)$ where $\bar{a}\in BV^{\star}([0,\hat{}\tau])$ allows for distributed feedback of $\eta^{\mathrm{P}}(z,t)$ as well as point evaluations for $z\in[0,\hat{}\tau)$.
+Similar to the SISO case, $\eta^{\mathrm{O}}_{1}(t)$ is a flat output of the system.
 II-C Benchmark system
 Consider a system of two vibrating strings
-$\partial_{t}^{2}x_{1}(z,t)-\partial_{z}^{2}x_{1}(z,t)$ $=0\quad z\in(-\alpha,0)$ (4a)
-$\partial_{t}^{2}x_{2}(z,t)-\partial_{z}^{2}x_{2}(z,t)$ $=0\quad z\in(0,\beta),$ (4b)
-coupled with a point mass at $z=0$ and thus subject to the dynamic boundary conditions
+$\partial_{t}^{2}x_{1}(z,t)-\partial_{z}^{2}x_{1}(z,t)$ $=0\quad z\in(-{l_{1}},0)$ (4a)
+$\partial_{t}^{2}x_{2}(z,t)-\partial_{z}^{2}x_{2}(z,t)$ $=0\quad z\in(0,{l_{2}}),$ (4b)
+coupled with a point mass M{M} at $z=0$ and thus subject to the dynamic boundary conditions
 $x_{1}(0,t)=x_{2}(0,t)$ $=\xi(t)$ (5a)
@@
 where $\xi(t)$ denotes the mass position as well as
-$\partial_{z}x_{1}(-\alpha,t)$ $=-u_{1}(t)$ (5c)
-$\partial_{z}x_{2}(\beta,t)$ $=u_{2}(t),$ (5d)
-with state [^2] In detail, $\bm{x}(\cdot,t)\!=\!(\xi(t),\dot{\xi}(t),x_{1}(\cdot,t),\partial_{t}x_{1}(\cdot,t),x_{2}(\cdot,t),\partial_{t}x_{2}(\cdot,t))^{\intercal}$ and $\mathcal{X}=\{\mathbb{R}^{2}\times L^{2}([-\alpha,0])\times H^{1}([-\alpha,0])\times L^{2}([0,\beta])\times H^{1}([0,\beta])\;|\;\xi(t)=x_{1}(0,t)=x_{2}(0,t)\}$.
+$\partial_{z}x_{1}(-{l_{1}},t)$ $=-u_{1}(t)$ (5c)
+$\partial_{z}x_{2}({l_{2}},t)$ $=u_{2}(t),$ (5d)
+with state [^2] In detail, $\bm{x}(\cdot,t)\!=\!(\xi(t),\dot{\xi}(t),x_{1}(\cdot,t),\partial_{t}x_{1}(\cdot,t),x_{2}(\cdot,t),\partial_{t}x_{2}(\cdot,t))^{\intercal}$ and $\mathcal{X}=\{\mathbb{R}^{2}\times H^{1}([-{l_{1}},0])\times L^{2}([-{l_{1}},0])\times H^{1}([0,{l_{2}}])\times L^{2}([0,{l_{2}}])\;|\;\xi(t)=x_{1}(0,t)=x_{2}(0,t)\}$.
 $\bm{x}(\cdot,t)\in\mathcal{X}$ as initially discussed in [18].
-By using the deflection and gradient at $z=0$ as Cauchy data, D’Alemberts’s formula yields the solutions
-$x_{i}(z,t)\!=\!\frac{1}{2}\!\left(x_{i}(0,t-z)\!+\!x_{i}(0,t+z)\right)\!+\!\frac{1}{2}\!\int\limits_{t-z}^{t+z}\partial_{z}x_{i}(0,\tau)\,\mathrm{d}\tau$ (6)
-for $i=1,2$ (c.f. [7, 4]), which reads
-$x_{i}(z,t)=C_{z}x_{i}(0,t)+\frac{1}{s}S_{z}\partial_{z}x_{i}(0,t)$ (7)
-in the algebraic setting, where $C_{z}\coloneqq\cosh(\partial_{t}z)=\frac{1}{2}(\sigma^{z}+\sigma^{-z})$ and $S_{z}\coloneqq\sinh(\partial_{t}z)=\frac{1}{2}(\sigma^{z}-\sigma^{-z})$.
-II-D Parametrisation
+II-D Input parametrisation by a flat output
 Starting at the interior mass, by using the tuple
 $\begin{pmatrix}{y}_{1}(t),{y}_{2}(t)\end{pmatrix}^{\intercal}=\begin{pmatrix}\xi(t),\partial_{z}x_{1}(0,t)+\partial_{z}x_{2}(0,t)\end{pmatrix}^{\intercal}$ (6)
-with $y_{1}\in H^{2}(\mathbb{R})$ and $y_{2}\in L^{2}(\mathbb{R})$ we directly obtain $\xi(t)=x_{1}(0,t)=x_{2}(0,t)={y}_{1}(t)$ as well as
-$\partial_{z}x_{1}(0,t)$ $=\frac{1}{2}\left({y}_{2}(t)-m\partial_{t}^{2}{y}_{1}(t)\right)$ (9a)
-$\partial_{z}x_{2}(0,t)$ $=\frac{1}{2}\left({y}_{2}(t)+m\partial_{t}^{2}{y}_{1}(t)\right)$ (9b)
-from (5b). With the Cauchy data known, (7) then yields the profiles, from which the input parameterisation
-$\begin{pmatrix}u_{1}(t)\\ u_{2}(t)\\ \end{pmatrix}=\begin{pmatrix}\frac{m}{2}\partial_{t}^{2}C_{\alpha}+\partial_{t}S_{\alpha}&-\frac{1}{2}C_{\alpha}\\ \frac{m}{2}\partial_{t}^{2}C_{\beta}+\partial_{t}S_{\beta}&\frac{1}{2}C_{\beta}\end{pmatrix}\begin{pmatrix}y_{1}(t)\\ y_{2}(t)\\ \end{pmatrix}$ (10)
+with $y_{1}\in H^{2}(\mathbb{R})$ as well as $y_{2}\in L^{2}(\mathbb{R})$ we directly obtain $\xi(t)=x_{1}(0,t)=x_{2}(0,t)={y}_{1}(t)$ and via (5b)
+$\partial_{z}x_{1}(0,t)$ $=\frac{1}{2}\left({y}_{2}(t)-{M}{\mathrm{d}_{t}}^{2}{y}_{1}(t)\right)$ (7a)
+$\partial_{z}x_{2}(0,t)$ $=\frac{1}{2}\left({y}_{2}(t)+{M}{\mathrm{d}_{t}}^{2}{y}_{1}(t)\right).$ (7b)
+By using the deflection $x_{i}(0,t)$ and its gradient $\partial_{z}x_{i}(0,t)$ as Cauchy data, the Method of Characteristics yields the profiles
+$x_{i}(z,t)\!=\!\frac{1}{2}\!\left(\!\xi(t-z)\!+\!\xi(t+z)\!+\!\!\int_{t-z}^{t+z}\!\partial_{z}x_{i}(0,\tau)\,\mathrm{d}\tau\right)$ (8)
+for $i=1,2$ (cf. [18, 5]), which can be written as
+$x_{i}(z,t)=C_{z}x_{i}(0,t)+S_{z}\partial_{z}x_{i}(0,t)$ (9)
+in the algebraic setting, where $C_{z}=\frac{1}{2}(\sigma^{z}+\sigma^{-z})$ and $S_{z}=\frac{1}{2}\frac{\sigma^{z}-\sigma^{-z}}{{\mathrm{d}_{t}}}$ employ $\sigma$ as well as $\mathrm{d}_{t}$} for shifts and derivatives.
+Herein, note that although $S_{z}$ contains a division by $\mathrm{d}_{t}$} which results in an integration in the parametrisation of $x_{i}(z,t)$ by $\bm{y}(t)$, this does not contradict the idea of a unique parametrisation as $S_{z}\in\mathcal{E}^{\prime}$ has compact support and the corresponding integral in (8) therefore is definite.
+With the parametrisation of the Cauchy data from (7), (9) then yields the profiles, thus, the input parametrisation
+$\begin{pmatrix}u_{1}(t)\\ u_{2}(t)\\ \end{pmatrix}=\begin{pmatrix}\frac{m}{2}{\mathrm{d}_{t}}^{2}C_{l_{1}}+{\mathrm{d}_{t}}S_{l_{1}}&-\frac{1}{2}C_{l_{1}}\\ \frac{m}{2}{\mathrm{d}_{t}}^{2}C_{l_{2}}+{\mathrm{d}_{t}}S_{l_{2}}&\frac{1}{2}C_{l_{2}}\end{pmatrix}\begin{pmatrix}y_{1}(t)\\ y_{2}(t)\\ \end{pmatrix}$ (10)
 is obtained via (5c) and (5d) showing that (6) is a flat output of the system (4)-(5).
 II-E Stating a controller form
-As known from, e.g [4], the problem to be solved for stating a controller form is to find a particular realisation of the input-output relation (10).
-A possible way to do so is to solve (10) for the highest derivatives of the largest predictions of $y_{1}(t)$ as well as $y_{2}(t)$ and to check for which initial conditions the resulting DDE system is well posed.
-However, taking the coefficients of the highest derivatives of the largest predictions in each column on the right hand-side of (10), the resulting leading column coefficient matrix
-$\operatorname{lccm}_{\rho,s}{\mathbf{H}}=\frac{1}{4}\begin{pmatrix}0&0\\ m&1\\ \end{pmatrix}$ (11)
-does not have full rank.
-This implies that (10) cannot directly be written as an explicit DDE system thus negating our approach.
-Therefore, more steps have to be taken for MIMO systems, which will be discussed in Section IV after we give a definition of the desired controller form for this case.
+As known from e.g.
+[5], the problem to be solved for stating a controller form is to find a particular realisation of the input-output relation (10).
+A possible way to do so is to solve (10) for the highest derivatives of the largest predictions of $y_{1}(t)$ as well as $y_{2}(t)$ and to check for which initial conditions the resulting FDE system is well-posed.
+However, the leading column coefficient matrix
+$\hat{}\mathbf{G}=\frac{1}{4}\begin{pmatrix}0&0\\ {M}&1\\ \end{pmatrix}$, (11)
+obtained by taking the coefficients of the highest derivatives of the largest predictions in each column on the right hand-side of (10), is singular.
+This implies that (10) cannot directly be written as an explicit FDE system thus negating our approach.
+Therefore, more steps must be taken in the MIMO case, which will be discussed in Section IV after we give a definition of a controller form for this system class.
 III A hyperbolic controller form
-In this section we define a general controller form for hyperbolic ODE-PDE MIMO systems and state several specific variants of the latter which may be helpful in different settings.
+In this section we define a general controller form for linear hyperbolic MIMO systems with dynamic boundary conditions and state several specific variants of the latter which may be helpful in different settings.
 III-A Definition
@@
 $\dot{\eta}^{\mathrm{O}}_{i,\nu_{i}-1}(t)$ $=\eta^{\mathrm{O}}_{i,\nu_{i}}(t)$ (12b)
-$\dot{\eta}^{\mathrm{O}}_{i,\nu_{i}}(t)$ $=\eta^{\mathrm{P}}_{i}(\check{\tau}_{i},t)$ (12c)
-$\partial_{t}\eta^{\mathrm{P}}_{i}(\tau,t)$ $=\partial_{\tau}\eta^{\mathrm{P}}_{i}(\tau,t),\quad\tau\in[\check{\tau}_{i},\hat{\tau}_{i})$ (12d)
-$\begin{split}\eta^{\mathrm{P}}_{i}(\hat{\tau}_{i},t)&=-\sum\limits_{j=1}^{m}\operatorname{\mathcal{S}}_{a^{\star}_{i,j}}\eta^{\mathrm{P}}_{j}(\cdot,t)+\bm{a}_{i,j}^{\intercal}\eta^{\mathrm{O}}_{j}(t)\\ &\qquad\qquad+\operatorname{\mathcal{S}}_{b_{i,j}^{\star}}\!*\,u_{j}(t),\end{split}$ (12e)
-with state $\bm{\eta}(\cdot,t)\in\mathbb{R}^{\nu_{\Sigma}}\times\prod_{i=1}^{m}L^{2}([\check{\tau}_{i},\hat{\tau}_{i}])$.
-Herein, the convolutions $\operatorname{\mathcal{S}}_{b_{i,j}^{\star}}\!*\,u_{j}(t)=\int_{\underline{\tau\!}}^{\bar{\tau}}u_{j}(t-\tau)\,\mathrm{d}b_{i,j}^{\star}(\tau)$ with the $b_{i,j}^{\star}$ to be further specified form the input map $\mathbf{B}^{\star}=[\operatorname{\mathcal{S}}_{b^{\star}_{i,j}}]$ which is invertible w.r.t.
+$\dot{\eta}^{\mathrm{O}}_{i,\nu_{i}}(t)$ $=\eta^{\mathrm{P}}_{i}(0,t)$ (12c)
+$\partial_{t}\eta^{\mathrm{P}}_{i}(\tau,t)$ $=\partial_{\tau}\eta^{\mathrm{P}}_{i}(\tau,t),\quad\tau\in[0,\hat{}\tau_{i})$ (12d)
+$\eta^{\mathrm{P}}_{i}(\hat{}\tau_{i},t)$ $=-\sum_{j=1}^{m}\Big(\left(a^{\star}_{ij}\eta^{\mathrm{P}}_{j}\right)(t)+\bm{a}_{ij}^{\intercal}\eta^{\mathrm{O}}_{j}(t)-\left(b_{ij}^{\star}u_{j}\right)(t)\Big)$ (12e)
+with state $\bm{\eta}(\cdot,t)\in\mathbb{R}^{\nu_{\Sigma}}\times\prod_{i=1}^{m}L^{2}([0,\hat{}\tau_{i}])$.
+Herein, the expressions $\left(b_{ij}^{\star}u_{j}\right)(t)=\int_{\underline{\tau}_{ij}\!}^{\bar{}\tau_{ij}}u_{j}(t+\tau)\,\mathrm{d}\bar{b}_{ij}(\tau)$ correspond to convolutions that – depending on the $\bar{b}_{ij}(\tau)$ – can represent input derivatives, predictions or delays.
+Furthermore, the input map $\mathbf{B}^{\star}=(b^{\star}_{ij})$ is invertible w.r.t.
 convolution.
@@
 However, the input of each block’s transport system now depends on the system input as well as the PDE and ODE states of all blocks.
-In the following, we will derive several special cases of the form (12) by specifying the functions $b_{i,j}^{\star}$ that define $\mathbf{B}^{\star}$.
+Note that similar to the integrator chains from the lumped MIMO case in section II-B2, there exist different ways to split the overall system delays into the $m$ transport systems.
+In the following, we will derive selected special cases of (12) by further specifying the functions $\bar{b}_{ij}(\tau)$.
 III-B A classic controller form
-To recover a classic form without any input derivatives, predictions or delays, we restrain the functions to $b_{i,j}^{\star}(t)=b_{i,j}H(t)$ with the Heavyside function $H(t)$, resulting in
-$\sum\limits_{j=1}^{m}\operatorname{\mathcal{S}}_{b^{\star}_{i,j}}\!*\,{u_{j}}(t)=\bm{b}_{i}^{\intercal}\bm{u}(t)$ (13)
-for $i\in\{1,\dotsc,{m}\}$ after evaluating the convolutions.
+To recover a classic form without input derivatives, predictions or delays in $\mathbf{B}^{\star}$, we restrain the functions to $\bar{b}_{ij}(\tau)=b_{ij}H(\tau)$ with the Heaviside step function $H(\tau)$, resulting in
+$\sum\limits_{j=1}^{m}\left(b^{\star}_{ij}u_{j}\right)(t)=\bm{b}_{i}^{\intercal}\bm{u}(t)$ (13)
+for $i\in\{1,\dotsc,{m}\}$ after evaluating the integrals.
 III-C Generalised controller forms
 III-C1 With input derivatives
-To allow for input derivatives to occur in (12e), we set $b_{i,j}^{\star}(t)=\sum_{k=0}^{D_{i,j}}b_{k}H^{(k)}(t)$, yielding
-$\sum\limits_{j=1}^{m}\operatorname{\mathcal{S}}_{b^{\star}_{i,j}}\!*\,{u_{j}}(t)=\sum\limits_{j=1}^{m}\sum_{k=0}^{D_{i,j}}b_{i,j,k}u_{j}^{(k)}(t)$ (14)
-where the $D_{i,j}$ denote the highest derivate order of $u_{j}(t)$ in the equations of subsystem $i$.
-Owing to the derivatives of the $u_{j}\in L^{2}(\mathbb{R}^{+})$, we propose the term discontinuous controller form for the subtype (14).
+To allow for input derivatives to occur in (12e), we set $\bar{b}_{ij}(\tau)=\sum_{k=0}^{D_{ij}}b_{ij,k}H^{(k)}(\tau)$, yielding
+$\sum\limits_{j=1}^{m}\left(b^{\star}_{ij}u_{j}\right)(t)=\sum\limits_{j=1}^{m}\sum_{k=0}^{D_{ij}}b_{ij,k}u_{j}^{(k)}(t)$ (14)
+where the $D_{i,j}$ denote the highest derivative order of $u_{j}(t)$ in the equations of subsystem $i$.
+Owing to the derivatives of the $u_{j}\in L^{2}_{\mathrm{loc}}(\mathbb{R}^{+})$, we propose the term discontinuous controller form for the subtype (14).
 III-C2 With input predictions
-For this case, we have $b_{i,j}^{\star}\in BV([-\bar{\tau}_{i,j},0])$, where the $\bar{\tau}_{i,j}$ denote the biggest predictions with which the input $u_{j}(t)$ appears in the subsystem $i$.
+For this case, we have $\bar{b}_{ij}\in\mathrm{BV}([0,\bar{}\tau_{ij}])$, where the $\bar{}\tau_{ij}$ denote the biggest predictions with which the input $u_{j}(t)$ appears in the subsystem $i$.
 As such a system depends on future inputs, we propose the term non-causal controller form for this case.
 III-C3 With input predictions and delays
-For a more general case, we allow $b_{i,j}^{\star}\in BV([-\bar{\tau}_{i,j},-\underline{\tau\!}_{i,j}])$, where the $\underline{\tau\!}_{i,j}$ denote the biggest delays of input $u_{j}(t)$.
+For a more general case, we allow $b_{ij}^{\star}\in\mathrm{BV}([\underline{\tau\!}_{ij},\bar{}\tau_{ij}])$, where the $\underline{\tau\!}_{ij}$ denote the biggest delays of input $u_{j}(t)$.
 As this form only yields a sensible state if the system is defined for $t\in\mathbb{R}$ instead of $\mathbb{R}^{+}$, we propose the term quasi controller form.
@@
 IV-A System class
-We discuss the linear first order hyperbolic MIMO system [^3] Note that the system from Section II-C can easily be transformed into this more general form which also allows for unsymmetrical predictions and delays in contrast to the more compact second order formulation.
+We discuss the linear first order hyperbolic MIMO system [^3] Note that the system from Section II-C can easily be transformed into this more general form which also allows for asymmetric predictions and delays in contrast to the more compact second order formulation.
 $\dot{\bm{\xi}}(t)$ $=\bm{F}\bm{\xi}(t)+\bm{B}\bm{x}^{-}(0,t)$ (15a)
@@
 $\bm{x}^{+}(0,t)$ $=\bm{Q}_{0}\bm{x}^{-}(0,t)+\bm{C}\bm{\xi}(t)$ (15c)
-$\bm{x}^{-}(1,t)$ $=\bm{Q}_{1}\bm{x}^{+}(1,t)+\bm{u}(t)$ (15d)
-as depicted in Figure 1, where $\bm{\xi}(t)\in\mathbb{R}^{n}$, $\bm{x}^{-}(\cdot,t)\in L^{2}([0,1])^{{n^{-}}}$ and $\bm{x}^{+}(\cdot,t)\in L^{2}([0,1])^{n^{+}}$ denote the boundary system’s as well as the backward and forward transport systems’ states.
+$\bm{x}^{-}(1,t)$ $=\bm{Q}_{1}\bm{x}^{+}(1,t)+\bm{u}(t),$ (15d)
+where $\bm{\xi}(t)\in\mathbb{R}^{n}$, $\bm{x}^{-}(\cdot,t)\in L^{2}([0,1],\mathbb{R}^{{n^{-}}})$ and $\bm{x}^{+}(\cdot,t)\in L^{2}([0,1],\mathbb{R}^{n^{+}})$ denote the boundary system’s as well as the backward and forward transport systems’ states.
 This model is also known as a PDE-ODE system in the backstepping literature, cf.
@@
 [19].
-For (15), we assume decoupled PDEs with $\bm{\Lambda}=\mathrm{diag}(\mathbf{\Lambda}^{-},-\mathbf{\Lambda}^{+})$ where $\mathbf{\Lambda}^{\pm}=\mathrm{diag}(\lambda_{1}^{\pm},\dotsc,\lambda_{n^{\pm}}^{\pm})$, $\lambda_{i}^{\pm}>\lambda_{i+1}^{\pm}>0$, the pair $(\bm{F},\bm{B})$ controllable and $\operatorname{rank}\bm{Q}_{0}={n^{+}}$ as well as $\operatorname{rank}\bm{Q}_{1}={n^{-}}$.
-Figure 1: Structure of the assumed system class
+For (15), we assume decoupled PDEs with $\bm{\Lambda}=\mathrm{diag}(\mathbf{\Lambda}^{-},-\mathbf{\Lambda}^{+})$ where $\mathbf{\Lambda}^{\pm}=\mathrm{diag}(\lambda_{1}^{\pm},\dotsc,\lambda_{n^{\pm}}^{\pm})$, $\lambda_{i}^{\pm}\geq\lambda_{i+1}^{\pm}>0$, the pair $(\bm{F},\bm{B})$ controllable and $\operatorname{rank}\bm{Q}_{0}=\operatorname{rank}\bm{Q}_{1}={n^{+}}$.
 IV-B Parametrisation
-As (15a) is controllable, a flatness-based parametrisation in the form of
-$\bm{\xi}(t)$ $=\mathbf{N}(\partial_{t})\bm{y}(t)$ (16a)
-$\bm{x}^{-}(0,t)$ $=\mathbf{D}(\partial_{t})\bm{y}(t)$ (16b)
+As (15a) is controllable, a flatness-based parametrisation
+$\bm{\xi}(t)$ $=\mathbf{N}({\mathrm{d}_{t}})\bm{y}(t)$ (16a)
+$\bm{x}^{-}(0,t)$ $=\mathbf{D}({\mathrm{d}_{t}})\bm{y}(t)$ (16b)
 can always be given.
-Furthermore, if the components of the flat output $\bm{y}(t)$ are chosen as the outputs of each block’s integrator chain of the systems’ MIMO controller form (2), the parametrisation matrix $\mathbf{D}(\partial_{t})$ is column reduced and $\operatorname{coldeg}_{i}\mathbf{N}(\partial_{t})<\operatorname{coldeg}_{i}\mathbf{D}(\partial_{t}),\;i\in\{1,\dotsc,m\}$ holds for the column degrees.
-Regarding the PDE subsystem, with $\bm{\xi}(t)$ as well as $\bm{x}^{-}(0,t)$ known, (15c) yields $\bm{x}^{+}(0,t)=\left(\mathbf{Q}_{0}\mathbf{D}(\partial_{t})+\mathbf{C}\mathbf{N}(\partial_{t})\right)\bm{y}(t)$ and thus the boundary values required to solve (15b) up to $z=1$.
+Furthermore, if the components of the flat output $\bm{y}(t)$ are chosen as the outputs of each block’s integrator chain of the systems’ MIMO controller form (2), the parametrisation matrix $\mathbf{D}({\mathrm{d}_{t}})$ is column reduced and $\operatorname{coldeg}_{i}\mathbf{N}({\mathrm{d}_{t}})<\operatorname{coldeg}_{i}\mathbf{D}({\mathrm{d}_{t}}),\;i\in\{1,\dotsc,m\}$ holds for the column degrees.
+Regarding the PDE subsystem, with $\bm{\xi}(t)$ as well as $\bm{x}^{-}(0,t)$ known, (15c) yields $\bm{x}^{+}(0,t)=\mathbf{\mathbf{H}}({\mathrm{d}_{t}})\bm{y}(t)=\left(\mathbf{Q}_{0}\mathbf{D}({\mathrm{d}_{t}})+\mathbf{C}\mathbf{N}({\mathrm{d}_{t}})\right)\bm{y}(t)$ and therefore the boundary values required to solve (15b) up to $z=1$.
 This yields
 $\bm{x}^{-}(1,t)$ $=\mathbf{P}(\sigma,t)\bm{x}^{-}(0,t)$ (17)
-$\bm{x}^{+}(1,t)$ $=\mathbf{\Delta}(\sigma,\partial_{t})\bm{x}^{+}(0,t)$ (18)
+$\bm{x}^{+}(1,t)$ $=\mathbf{\Delta}(\sigma,{\mathrm{d}_{t}})\bm{x}^{+}(0,t)$ (18)
 with the prediction matrix $\mathbf{P}(\sigma)=\operatorname{diag}(\sigma^{\tau_{1}^{-}},\dotsc,\sigma^{\tau_{{n^{-}}}^{-}})$ and the delay matrix $\mathbf{\Delta}(\sigma)=\operatorname{diag}(\sigma^{-\tau_{1}^{+}},\dotsc,\sigma^{-\tau_{{n^{+}}}^{+}})$ where $\tau_{i}^{\pm}=\frac{1}{\lambda_{i}^{\pm}}$.
 Finally, combining Equations 16a, 16b, 17 and 18 with (15d) yields the matrix
-$\mathbf{H}(\sigma,\partial_{t})\!=\!\mathbf{P}(\sigma)\mathbf{D}(\partial_{t})-\mathbf{Q}_{1}\mathbf{\Delta}(\sigma)\left(\mathbf{Q}_{0}\mathbf{D}(\partial_{t})+\mathbf{C}\mathbf{N}(\partial_{t})\right)$ (19)
+$\mathbf{G}(\sigma,{\mathrm{d}_{t}})\!=\!\mathbf{P}(\sigma)\mathbf{D}({\mathrm{d}_{t}})-\mathbf{Q}_{1}\mathbf{\Delta}(\sigma)\mathbf{H}({\mathrm{d}_{t}})$ (19)
 for the parametrisation
-$\bm{u}(t)=\mathbf{H}(\sigma,\partial_{t})\bm{y}(t)$ (20)
+$\bm{u}(t)=\mathbf{G}(\sigma,{\mathrm{d}_{t}})\bm{y}(t)$ (20)
 of the system input by the flat output of the ODE.
-As was already to be expected from the motivational example’s parametrisation (10), (19) now structurally shows that $\mathbf{\mathbf{P}}(\sigma)$ and $\mathbf{\Delta}(\sigma)$ introduce all predictions and delays that occur in the system into every column of $\mathbf{H}(\sigma,\partial_{t})$ which poses a problem.
-Next, we will extend the notion of polynomials in $\partial_{t}$ to incorporate shifts and remedy this inconvenience.
-IV-C Generalised polynomials
-With the set $T=\{\tau_{1}^{-},\dotsc,\tau_{n^{-}}^{-},\tau_{1}^{+},\dotsc,\tau_{n^{+}}^{+},\}$ of all unique predictions and shifts in the system, by interpreting the members of $\Sigma=\{\sigma^{\tau},\sigma^{-\tau}|\tau\in T\}$ and $\partial_{t}$ as independent generators, we conceive the entries $h_{i,j}(\sigma,\partial_{t})$ of $\mathbf{H}(\sigma,\partial_{t})$ to be from $\mathcal{R}_{\mathrm{p}}=\mathbb{R}[\Sigma,\partial_{t}]$, the space of generalised polynomials in $\sigma$ with real exponents, whose coefficients are polynomials in $\partial_{t}$ with coefficients from $\mathbb{R}$.
+The structure of (19) now shows that $\mathbf{\mathbf{P}}(\sigma)$ and $\mathbf{\Delta}(\sigma)$ introduce all predictions and delays of the system into every column of $\mathbf{G}(\sigma,{\mathrm{d}_{t}})$ which is the root cause of the problem with the motivational example’s parametrisation (10).
+Next, we will employ quasipolynomials to resolve the issue.
+IV-C Quasipolynomials
+With the set $T=\{\tau_{1}^{-},\dotsc,\tau_{n^{-}}^{-},\tau_{1}^{+},\dotsc,\tau_{n^{+}}^{+},\}$ of all unique predictions and shifts in the system, by interpreting the members of $\Sigma=\{\sigma^{\tau},\sigma^{-\tau}|\tau\in T\}$ and $\mathrm{d}_{t}$} as independent generators, we conceive the entries $g_{i,j}(\sigma,{\mathrm{d}_{t}})$ of $\mathbf{G}(\sigma,{\mathrm{d}_{t}})$ to be from $\mathcal{R}_{\mathrm{p}}=\mathbb{R}[\Sigma,{\mathrm{d}_{t}}]$, the ring of quasipolynomials in $\sigma$ with real exponents, whose coefficients are polynomials in $\mathrm{d}_{t}$} with coefficients from $\mathbb{R}$.
 Furthermore, for the elements
-$r(\sigma,\partial_{t})=\sum_{i=1}^{n}a_{i}(\partial_{t})\sigma^{\alpha_{i}}\in\mathcal{R}_{\mathrm{p}}$ (21)
-with $a_{i}(\partial_{t})\in\mathbb{R}[\partial_{t}]$, the ordering $\alpha_{i}>\alpha_{i+1}$, for $i\in\{1,\dotsc,n-1\}$ is assumed.
+$r(\sigma,{\mathrm{d}_{t}})=\sum_{i=1}^{n}a_{i}({\mathrm{d}_{t}})\sigma^{\alpha_{i}}\in\mathcal{R}_{\mathrm{p}}$ (21)
+with $a_{i}({\mathrm{d}_{t}})\in\mathbb{R}[{\mathrm{d}_{t}}]$, the ordering $\alpha_{i}>\alpha_{i+1}$, for $i\in\{1,\dotsc,n-1\}$ is assumed.
 To ease working on the polynomials (21), we define the pseudo degrees
-$\deg_{\sigma}^{+}r(\sigma,\partial_{t})\coloneq\max_{i\in\{1,\dotsc,n\}}\alpha_{i},\;\deg_{\sigma}^{-}r(\sigma,\partial_{t})\coloneq\min_{i\in\{1,\dotsc,n\}}\alpha_{i}$
-as well as the degree [^4] This choice is motivated by the case of Laurent polynomials, which posses positive and negative integer exponents and form an euclidean ring for the given degree definition, cf.
+$\deg_{\sigma}^{+}r(\sigma,{\mathrm{d}_{t}})\coloneq\max_{i\in\{1,\dotsc,n\}}\alpha_{i},\;\deg_{\sigma}^{-}r(\sigma,{\mathrm{d}_{t}})\coloneq\min_{i\in\{1,\dotsc,n\}}\alpha_{i}$
+as well as the degree [^4] This choice is motivated by the case of Laurent polynomials, which possess positive and negative integer exponents and form an euclidean ring for the given degree definition, cf.
 [10, 20].
-$\deg_{\sigma}r(\sigma,\partial_{t})\coloneq\deg_{\sigma}^{+}r(\sigma,\partial_{t})-\deg_{\sigma}^{-}r(\sigma,\partial_{t}),$ (22)
+$\deg_{\sigma}r(\sigma,{\mathrm{d}_{t}})\coloneq\deg_{\sigma}^{+}r(\sigma,{\mathrm{d}_{t}})-\deg_{\sigma}^{-}r(\sigma,{\mathrm{d}_{t}}),$ (22)
 yielding the upper and lower boundary as well as the diameter of the interval, respectively. Next, we discuss how these intervals can be reduced in a systematic way.
-IV-D Shift interval reduction
-IV-D1 Preliminary Transformations
+IV-D A shift interval reduction algorithm
+After stating the requirements on the input parametrisation in terms of degree conditions for quasipolynomials, this section shows how the shifts in single entries of $\mathbf{G}(\sigma,{\mathrm{d}_{t}})$ can be reduced before the element-wise operations are condensed into an algorithm to reduce the shifts in the whole matrix.
+IV-D1 Requirements
+To obtain the controller form (12), the input parametrisation (20) must be solved for the highest derivatives of the flat output.
+Furthermore, to introduce a HCF state in a direct fashion, these must occur together with the largest predictions as well as largest delays and the total system shift $\tau_{\Sigma}=\sum_{i=1}^{{n^{-}}}\tau^{-}_{i}+\sum_{i=1}^{{n^{+}}}\tau^{+}_{i}$ must be preserved.
+In the algebraic setting, this corresponds to the conditions
+$\operatorname{lccm}_{{\mathrm{d}_{t}}}(\operatorname{lccm}^{+}_{\sigma}\mathbf{G}(\sigma,{\mathrm{d}_{t}}))$ $=\operatorname{lccm}^{+}_{\sigma}(\operatorname{lccm}_{{\mathrm{d}_{t}}}\mathbf{G}(\sigma,{\mathrm{d}_{t}}))$ (23a)
+$\sum_{i=1}^{m}\operatorname{coldeg}_{\sigma,i}\mathbf{G}(\sigma,{\mathrm{d}_{t}})$ $=\tau_{\Sigma}$ (23b)
+where (23a) must have full rank.
+In this setting, the operator $\operatorname{lccm}^{+}_{\sigma}$ denotes the LCCM w.r.t.
+$\sigma$ while using the pseudo degree $\deg^{+}_{\sigma}$ to identify the leading coefficients.
+Thus, (23a) is trivially satisfied if the degree conditions
+$\deg^{+}_{\sigma}{g}_{ij}(\sigma,{\mathrm{d}_{t}})$ $\leq\deg^{+}_{\sigma}{g}_{jj}(\sigma,{\mathrm{d}_{t}})$ (24a)
+$\deg_{{\mathrm{d}_{t}}}{g}_{ij}(\sigma,{\mathrm{d}_{t}})$ $\leq\deg_{{\mathrm{d}_{t}}}{g}_{jj}(\sigma,{\mathrm{d}_{t}})$ (24b)
+hold for $i,j\in\{1,\dotsc,{m}\}$ as all LCCMs then are diagonal.
+IV-D2 Preliminary transformations
 Lemma 1
-By permuting the input and/or flat output, $\mathbf{H}(\sigma,\partial_{t})$ can be sorted, such that the degrees w.r.t.
-$\sigma$ in each column are ascending and the highest degrees w.r.t.
-$\partial_{t}$ appear on the main diagonal, thus with $j\in\{1,\dotsc,m\}$
-$\deg_{\sigma}{h}_{i,j}(\sigma,\partial_{t})$ $\leq\deg_{\sigma}{h}_{i+1,j}(\sigma,\partial_{t}),$ (23a)
-$\deg_{\partial_{t}}{h}_{l,j\neq}(\sigma,\partial_{t})$ $\leq\deg_{\partial_{t}}{h}_{l,l}(\sigma,\partial_{t})$ (23b)
-hold for $i\in\{1,\dotsc,m-1\}$ and $l\in\{1,\dotsc,m\}$.
+By transforming the system input and permuting the components of the flat output, a matrix $\mathbf{\tilde{G}}(\sigma,{\mathrm{d}_{t}})$ can be obtained, such that a) the leading row delays $d_{i}=\operatorname{rowdeg}_{\sigma,i}^{-}\mathbf{\tilde{G}}(\sigma,{\mathrm{d}_{t}})$ coincide, b) the leading row degrees $r_{i}=\operatorname{rowdeg}_{\sigma,i}\mathbf{\tilde{G}}(\sigma,{\mathrm{d}_{t}})$ are ascending and c) the highest degrees w.r.t.
+$\mathrm{d}_{t}$} in each column are found on the main diagonal.
+Thus, $\mathbf{\tilde{G}}(\sigma,{\mathrm{d}_{t}})$ satisfies (24b) as well as
+$d_{1}$ $=d_{2}=\ldots=d_{m},$ (25a)
+$r_{i}$ $\leq r_{i+1},\;i\in\{1,\dotsc,m-1\}.$ (25b)
 Proof:
-As from (19), there are equal predictions and delays in each row of $\mathbf{H}(\sigma,\partial_{t})$ due to the diagonal structure of $\mathbf{P}(\sigma)$ and $\mathbf{\Delta}(\sigma)$.
-Then, by (22) the same follows for their corresponding degrees which is why the ordering (23a) is easily obtained by permuting the flat output.
-Now, from the parametrisation (16) we have that $\operatorname{coldeg}\mathbf{N}(\partial_{t})<\operatorname{coldeg}\mathbf{D}(\partial_{t})$, which is why $\mathbf{C}$ has no effect regarding the degrees of $\partial_{t}$} in (19) and also that $\mathbf{D}(\partial_{t})$ is column reduced.
-Since this property is preserved under row transformations over $\mathbb{R}$ and $\mathbf{Q}_{0},\mathbf{Q}_{1}$ have full rank, we obtain that $\mathbf{H}(\sigma,\partial_{t})$ is column reduced w.r.t.
-$\partial_{t}$.
-Finally, as the elements in each row have equal degree w.r.t.
-$\sigma$, (23b) is achieved by permuting the input.
+Due to the structure of (19), a fully populated $\mathbf{Q}_{1}$ introduces larger delays from lower rows into rows further up.
+Therefore, we use a lower left triangular row transform (obtainable e.g.
+from the LU decomposition of $\mathbf{Q}_{1}$) to bring $\mathbf{Q}_{1}$ into upper left triangular form.
+This eliminates superfluous delays without increasing the predictions as $\mathbf{P}(\sigma)$ is diagonal.
+Next, the leading row delays are harmonised at their mean $\bar{d}=\sum_{i=1}^{m}d_{i}$ by left multiplying with the unimodular $\mathbf{S}(\sigma)=\operatorname{diag}(\sigma^{\bar{d}-d_{1}},\dotsc,\sigma^{\bar{d}-d_{m}})$.
+Finally, (25b) is obtained by permuting the input components.
+Furthermore, from the parametrisation (16) we have that $\operatorname{coldeg}_{i}\mathbf{N}({\mathrm{d}_{t}})<\operatorname{coldeg}_{i}\mathbf{D}({\mathrm{d}_{t}})$ (which is why $\mathbf{C}$ has no effect regarding the degrees in (19) w.r.t.
+$\mathrm{d}_{t}$}}) and also that $\mathbf{D}({\mathrm{d}_{t}})$ is column reduced.
+Since this property is preserved under transformations over $\mathbb{R}$ and $\mathbf{Q}_{0},\mathbf{Q}_{1}$ have full rank, we obtain that $\mathbf{G}(\sigma,{\mathrm{d}_{t}})$ is column reduced w.r.t.
+$\mathrm{d}_{t}$}.
+Therefore, (24b) is achieved by permuting the components of $\bm{y}(t)$.
 ∎
-IV-D2 Goal
-As discussed in the motivation, to obtain the controller form (12), the input parametrisation (20) must be solved for the highest derivatives of the flat output.
-Furthermore, to introduce a HCF state in a direct fashion, these must occur together with the largest predictions as well as largest delays.
-In the algebraic setting, this corresponds to the composition of taking the LCCMs w.r.t.
-$\sigma$ and $\partial_{t}$ of $\mathbf{H}(\sigma,\partial_{t})$ to commute:
-$\operatorname{lccm}_{s}(\operatorname{lccm}^{+}_{\sigma}\mathbf{H}(\sigma,\partial_{t}))$ $=\operatorname{lccm}^{+}_{\sigma}(\operatorname{lccm}_{s}\mathbf{H}(\sigma,\partial_{t}))$ (24a)
-$\operatorname{lccm}_{s}(\operatorname{lccm}_{\sigma}\mathbf{H}(\sigma,\partial_{t}))$ $=\operatorname{lccm}_{\sigma}(\operatorname{lccm}_{s}\mathbf{H}(\sigma,\partial_{t}))$ (24b)
-and the result having full rank. The condition for $\deg_{\sigma}^{-}$ then follows from (22). These requirements are trivially satisfied if the degree conditions
-$\deg^{+}_{\sigma}{h}_{i,j\neq i}(\sigma,\partial_{t})$ $\leq\deg^{+}_{\sigma}{h}_{i,i}(\sigma,\partial_{t})$ (25a)
-$\deg_{\sigma}{h}_{i,j\neq i}(\sigma,\partial_{t})$ $\leq\deg_{\sigma}{h}_{i,i}(\sigma,\partial_{t})$ (25b)
-$\deg_{\partial_{t}}{h}_{i,j\neq i}(\sigma,\partial_{t})$ $\leq\deg_{\partial_{t}}{h}_{i,i}(\sigma,\partial_{t})$ (25c)
-hold for all $i,j\in\{1,\dotsc,{m}\}$, since all LCCMs in this case are diagonal.
-Furthermore, as by Lemma 1, (25a),(25b) for $j\leq i$ as well as (25c) for $j\in\{1,\dotsc,m\}$ are satisfied, the elements $h_{i,i}(\sigma,\partial_{t})$ of each column are now used to reduce the (pseudo) degrees of the entries below them, such that all
-$\bar{h}_{j,i}(\sigma,\partial_{t})=h_{i,j}(\sigma,\partial_{t})-q_{j,i}(\sigma,\partial_{t})h_{i,i}(\sigma,\partial_{t})$ (26)
-for $i,j\in\{1,\dotsc,m\},j>i$ satisfy (25). Next, we discuss a generalised polynomial long division (GPLD) algorithm to perform such a reduction for a single element.
-IV-D3 Generalised polynomial long division
-For now, we treat all entries in (26) as elements of $\mathcal{R}_{\mathrm{r}}=\mathbb{R}(\partial_{t})\left[\Sigma\right]$ which allows for rational expressions in $\partial_{t}$ as coefficients and therefore enables us to develop a polynomial long division scheme regarding $\sigma$.
+In $\mathbf{\tilde{G}}(\sigma,{\mathrm{d}_{t}})$ from Lemma 1, the diagonal elements $g_{ii}(\sigma,{\mathrm{d}_{t}})$ of each column can be used to reduce the (pseudo) degrees of the entries below them, such that all
+$\bar{g}_{ij}(\sigma,{\mathrm{d}_{t}})=\tilde{g}_{ij}(\sigma,{\mathrm{d}_{t}})-q_{ij}(\sigma,{\mathrm{d}_{t}})\tilde{g}_{jj}(\sigma,{\mathrm{d}_{t}})$ (26)
+for $i,j\in\{1,\dotsc,m\},i>j$ satisfy (24).
+Next, we discuss how to compute the required $q_{ij}(\sigma,{\mathrm{d}_{t}})$ to perform such a reduction and develop a modified quasipolynomial long division (QPLD) algorithm.
+IV-D3 Quasipolynomial long division on $\mathcal{R}_{\mathrm{r}}$
+Treating all entries in (26) as elements of $\mathcal{R}_{\mathrm{r}}=\mathbb{R}({\mathrm{d}_{t}})\left[\Sigma\right]$ which allows for rational expressions in $\mathrm{d}_{t}$} as coefficients enables us to develop a tentative polynomial long division scheme in $\sigma$:
 Lemma 2
-Let $x(\sigma,\partial_{t}),y(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{r}},\;y(\sigma,\partial_{t})\neq 0$ be generalised polynomials with rational coefficients where $\deg_{\sigma}x(\sigma)\geq\deg_{\sigma}y(\sigma)$.
-Then there exist $q(\sigma,\partial_{t}),r(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{r}}$ such that
-$x(\sigma,\partial_{t})=q(\sigma,\partial_{t})y(\sigma,\partial_{t})+r(\sigma,s)$ (27)
-with $\deg_{\sigma}r(\sigma,\partial_{t})<\deg_{\sigma}y(\sigma,\partial_{t})$ as well as $\deg^{+}_{\sigma}r(\sigma,\partial_{t})<\deg^{+}_{\sigma}y(\sigma,\partial_{t})$ and $\deg^{-}_{\sigma}r(\sigma,\partial_{t})>\deg^{-}_{\sigma}y(\sigma,\partial_{t})$.
+Let $x(\sigma,{\mathrm{d}_{t}}),y(\sigma,{\mathrm{d}_{t}})\in\mathcal{R}_{\mathrm{r}},\;y(\sigma,{\mathrm{d}_{t}})\neq 0$ be quasipolynomials in $\sigma$ with rational coefficients in $\mathrm{d}_{t}$} and $\deg_{\sigma}x(\sigma)\geq\deg_{\sigma}y(\sigma)$ as well as $\deg_{{\mathrm{d}_{t}}}x(\sigma)\leq\deg_{{\mathrm{d}_{t}}}y(\sigma)$.
+Then there exist $q(\sigma,{\mathrm{d}_{t}}),r(\sigma,{\mathrm{d}_{t}})\in\mathcal{R}_{\mathrm{r}}$ such that
+$x(\sigma,{\mathrm{d}_{t}})=q(\sigma,{\mathrm{d}_{t}})y(\sigma,{\mathrm{d}_{t}})+r(\sigma,s)$ (27)
+with $\deg_{\sigma}r(\sigma,{\mathrm{d}_{t}})<\deg_{\sigma}y(\sigma,{\mathrm{d}_{t}})$, $\deg^{+}_{\sigma}r(\sigma,{\mathrm{d}_{t}})<\deg^{+}_{\sigma}y(\sigma,{\mathrm{d}_{t}})$ and $\deg^{-}_{\sigma}r(\sigma,{\mathrm{d}_{t}})>\deg^{-}_{\sigma}y(\sigma,{\mathrm{d}_{t}})$.
 Proof:
-As the following steps are valid for arbitrary coefficients, the explicit mention of $\partial_{t}$ will be omitted.
+As the following steps are valid for arbitrary coefficients, the explicit mention of $\mathrm{d}_{t}$} will be omitted.
 In the general case we have $\deg_{\sigma}x(\sigma)>\deg_{\sigma}y(\sigma)$ as well as $\deg^{+}_{\sigma}x(\sigma)>\deg^{+}_{\sigma}y(\sigma)$, $\deg^{-}_{\sigma}x(\sigma)<\deg^{-}_{\sigma}y(\sigma)$ or both.
-Since the exponents of the polynomials $x(\sigma,\partial_{t})=\sum_{i=1}^{n}a_{i}(\partial_{t})\sigma^{\alpha_{i}}$, $y(\sigma,\partial_{t})=\sum_{i=1}^{m}b_{i}(\partial_{t})\sigma^{\beta_{i}}$ are ordered, we use
-$q(\sigma)=\begin{cases}\frac{a_{1}\sigma^{\alpha_{1}}}{b_{1}\sigma^{\beta_{1}}},&\text{ if }\deg_{\sigma}^{+}x(\sigma)>\deg^{+}_{\sigma}y(\sigma)\\ \frac{a_{n}\sigma^{\alpha_{n}}}{b_{m}\sigma^{\beta_{m}}},&\text{ if }\deg^{-}_{\sigma}x(\sigma)<\deg^{-}_{\sigma}y(\sigma),\end{cases}$ (28)
+Since the exponents of $x(\sigma,{\mathrm{d}_{t}})=\sum_{i=1}^{u}a_{i}({\mathrm{d}_{t}})\sigma^{\alpha_{i}}$, $y(\sigma,{\mathrm{d}_{t}})=\sum_{i=1}^{v}b_{i}({\mathrm{d}_{t}})\sigma^{\beta_{i}}$ are ordered, we use
+$q(\sigma)=\begin{cases}\frac{a_{1}}{b_{1}}\sigma^{\alpha_{1}-\beta_{1}},&\text{ if }\deg_{\sigma}^{+}x(\sigma)>\deg^{+}_{\sigma}y(\sigma)\\ \frac{a_{u}}{b_{v}}\sigma^{\alpha_{u}-\beta_{v}},&\text{ if }\deg^{-}_{\sigma}x(\sigma)<\deg^{-}_{\sigma}y(\sigma),\end{cases}$ (28)
 where in the case that both are true, the upper branch is taken. Assuming the upper branch was taken, this yields
-$\begin{split}r(\sigma)=x(\sigma)-q(\sigma)y(\sigma)=\sum\limits_{i=2}^{n}a_{i}\sigma^{\alpha_{i}}-\frac{a_{1}\sigma^{\alpha_{1}}}{b_{1}\sigma^{\beta_{1}}}\sum\limits_{i=2}^{m}b_{i}\sigma^{\beta_{i}},\end{split}$
+$\begin{split}r(\sigma)=x(\sigma)-q(\sigma)y(\sigma)=\sum\limits_{i=2}^{u}a_{i}\sigma^{\alpha_{i}}\!-\!\frac{a_{1}}{b_{1}}\sigma^{\alpha_{1}-\beta_{1}}\!\sum\limits_{i=2}^{v}b_{i}\sigma^{\beta_{i}}\end{split}$
 for which we find that
 $\deg^{+}_{\sigma}r(\sigma)=\max\left\{\alpha_{2},\alpha_{1}-(\beta_{1}-\beta_{2})\right\}<\deg^{+}_{\sigma}x(\sigma)$
-as well as $\deg^{-}_{\sigma}r(\sigma)=\min\{\alpha_{n},\alpha_{1}-(\beta_{1}-\beta_{m})\}$ in which we identify $\deg_{\sigma}y(\sigma)=\beta_{1}-\beta_{m}$ and use $\alpha_{1}=\deg_{\sigma}x(\sigma)+\alpha_{n}$ as well as $\kappa=\deg_{\sigma}x(\sigma)-\deg_{\sigma}y(\sigma)\geq 0$ to obtain
-$\begin{split}\deg^{-}_{\sigma}r(\sigma)=\min\{&\alpha_{n},\alpha_{n}+\kappa\}=\alpha_{n}=\deg^{-}_{\sigma}x(\sigma).\end{split}$
+as well as $\deg^{-}_{\sigma}r(\sigma)=\min\{\alpha_{u},\alpha_{1}-(\beta_{1}-\beta_{v})\}$ in which we identify $\deg_{\sigma}y(\sigma)=\beta_{1}-\beta_{v}$ and use $\alpha_{1}=\deg_{\sigma}x(\sigma)+\alpha_{u}$ as well as $\kappa=\deg_{\sigma}x(\sigma)-\deg_{\sigma}y(\sigma)\geq 0$ to obtain
+$\begin{split}\deg^{-}_{\sigma}r(\sigma)=\min\{&\alpha_{u},\alpha_{u}+\kappa\}=\alpha_{u}=\deg^{-}_{\sigma}x(\sigma).\end{split}$
 This shows that the predictions in $x(\sigma)$ can be reduced without affecting the delays.
-Regarding the overall degree, we have $\deg_{\sigma}r(\sigma)=\max\{\alpha_{2}-\alpha_{n},\alpha_{1}-\alpha_{n}-(\beta_{1}-\beta_{2})\}$ which yields $\deg_{\sigma}r(\sigma)<\deg_{\sigma}x(\sigma)$.
+Regarding the overall degree, we have $\deg_{\sigma}r(\sigma)=\max\{\alpha_{2}-\alpha_{u},\alpha_{1}-\alpha_{u}-(\beta_{1}-\beta_{2})\}$ which yields $\deg_{\sigma}r(\sigma)<\deg_{\sigma}x(\sigma)$.
 Hence, $x(\sigma)$ can be reduced in a manner such that the iteration
 $r_{k}(\sigma)=r_{k-1}(\sigma)-q_{k}(\sigma)y(\sigma)$
-with $r_{0}(\sigma)=x(\sigma)$ and $q_{k}(\sigma)=\frac{c_{k,1}\sigma^{\gamma_{k,1}}}{b_{1}\sigma^{\beta_{1}}}$ by $r_{k}(\sigma)=\sum_{i=1}^{n}c_{k,i}\sigma^{\gamma_{k,i}}$ will for some $k^{+}$ terminate with $\deg^{+}_{\sigma}r_{k^{+}}(\sigma)<\deg^{+}_{\sigma}y(\sigma)$ since $x(\sigma)$ only has a finite amount of summands.
-Now, (if required) the second branch of (28) is taken and similar computations yield that restarting the iteration with $r_{0}(\sigma)=r_{k^{+}}(\sigma)$ and choosing $q_{k}(\sigma)=\frac{c_{k,1}\sigma^{\gamma_{k,1}}}{b_{m}\sigma^{\beta_{m}}}$ will yield $\deg^{-}_{\sigma}r_{k^{-}}(\sigma)>\deg^{-}_{\sigma}y(\sigma)$ after $k^{-}$ more steps.
-Summing up all $q_{k}(\sigma)$ from both iterations then yields the resulting “quotient” $q(\sigma)$ with common denominator $d=b_{1}^{k^{+}}b_{m}^{k^{-}}$.
+with $r_{0}(\sigma)=x(\sigma)$ and $q_{k}(\sigma)=\frac{c_{k,1}}{b_{1}}\sigma^{\gamma_{k,1}-\beta_{1}}$ by $r_{k}(\sigma)=\sum_{i=1}^{w_{k}}c_{k,i}\sigma^{\gamma_{k,i}}$ will for some $k^{+}$ terminate with $\deg^{+}_{\sigma}r_{k^{+}}(\sigma)<\deg^{+}_{\sigma}y(\sigma)$ since $x(\sigma)$ only has a finite number of summands.
+If required, the second branch of (28) is then taken and similar computations yield that restarting the iteration with $r_{0}(\sigma)=r_{k^{+}}(\sigma)$ and choosing $q_{k}(\sigma)=\frac{c_{k,1}}{b_{v}}\sigma^{\gamma_{k,1}-\beta_{v}}$ will yield $\deg^{-}_{\sigma}r_{k^{-}}(\sigma)>\deg^{-}_{\sigma}y(\sigma)$ after $k^{-}$ more steps.
+Summing up all $q_{k}(\sigma)$ from both iterations then yields the resulting “quotient” $q(\sigma)$ with denominator $d=b_{1}^{k^{+}}b_{v}^{k^{-}}$.
 ∎
-With Lemma 2, the shift degrees are taken care of.
-However, the coefficients of $q(\sigma,\partial_{t}),r(\sigma,\partial_{t})$ are rational expressions in $\partial_{t}$.
-Furthermore, in the analytic setting, their denominators can be interpreted as functions that their numerator polynomials are applied to.
-This results in piecewise continuous functions which do not necessarily have compact support.
-In other words: They generally do not correspond to differential or (distributed) delay operators with finite shift amplitude from $\mathcal{E}^{\prime}$.
-However, thanks to the Palay-Wiener theorem, if their Laplace transforms are entire functions, they indeed are such operators, cf.
-[11, Thm.
-7.3.1].
+IV-D4 Quasipolynomial long division on $\mathcal{R}_{\mathrm{e}}$
+With Lemma 2, the shift degrees have been taken care of.
+However, the coefficients of $q(\sigma,{\mathrm{d}_{t}})$ and $r(\sigma,{\mathrm{d}_{t}})$ still are rational expressions in $\mathrm{d}_{t}$}.
+Since, in the analytic setting, their denominators can be interpreted as functions that their numerator polynomials are applied to, the results are piecewise continuous functions which do not necessarily have compact support.
+In other words: They generally do not correspond to differential or (distributed) delay operators with finite shift amplitude from $\mathcal{E}^{\prime}$ and can therefore not be used to reduce the shifts as they would introduce indefinite integrals.
+Nevertheless, thanks to the Paley-Wiener theorem [21, Thm.
+7.3.1], if their Laplace transforms are entire functions, they indeed are such operators.
+For example, take $S_{z}=\frac{1}{2}\frac{\sigma^{z}-\sigma^{-z}}{{\mathrm{d}_{t}}}\in\mathcal{R}_{\mathrm{r}}$ from the introduction, whose Laplace transform $\mathcal{L}\{S_{z}\}(s)=\frac{1}{2s}\left(e^{sz}-e^{-sz}\right)$ appears to possess a pole at $s=0$.
+However, since its numerator introduces a zero at the same location, this pole can be lifted rendering $\mathcal{L}\{S_{z}\}(s)$ an entire function and thus yielding $S_{z}\in\mathcal{E}^{\prime}$ (which was already clear from (8)).
+Thus, the idea is to introduce numerator zeros to $q(\sigma,{\mathrm{d}_{t}})$ in form of a correction term $p({\mathrm{d}_{t}})$ such that a modified $q^{\star}(\sigma,{\mathrm{d}_{t}})=q(\sigma,{\mathrm{d}_{t}})+p({\mathrm{d}_{t}})$ belongs to $\mathcal{R}_{\mathrm{e}}=\mathcal{R}_{\mathrm{r}}\cap\mathcal{E}^{\prime}$ and is therefore admissible.
 Lemma 3
-Let $x(\sigma,\partial_{t}),y(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{p}},\;y(\sigma,\partial_{t})\neq 0$ be generalised polynomials with polynomial coefficients where $\deg_{\sigma}x(\sigma)\geq\deg_{\sigma}y(\sigma)$, as well as $\deg_{\partial_{t}}x(\sigma)\leq\deg_{\partial_{t}}y(\sigma)$.
-Then there exist $q^{\star}(\sigma,\partial_{t}),r^{\star}(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{w}}=\mathcal{R}_{\mathrm{r}}\cap\mathcal{E}^{\prime}$ such that
-$x(\sigma,\partial_{t})=q^{\star}(\sigma,\partial_{t})y(\sigma,\partial_{t})+r^{\star}(\sigma,s)$ (29)
-with $\deg_{\sigma}r^{\star}(\sigma,\partial_{t})=\deg_{\sigma}y(\sigma,\partial_{t})$, $\deg^{+}_{\sigma}r^{\star}(\sigma,\partial_{t})=\deg^{+}_{\sigma}y(\sigma,\partial_{t})$ and $\deg^{-}_{\sigma}r^{\star}(\sigma,\partial_{t})=\deg^{-}_{\sigma}y(\sigma,\partial_{t})$.
+Let $x(\sigma,{\mathrm{d}_{t}}),y(\sigma,{\mathrm{d}_{t}})\in\mathcal{R}_{\mathrm{e}},\;y(\sigma,{\mathrm{d}_{t}})\neq 0$ be operators from $\mathcal{E}^{\prime}$, written as quasipolynomials in $\sigma$ with rational coefficients in $\mathrm{d}_{t}$} and $\deg_{\sigma}x(\sigma)\geq\deg_{\sigma}y(\sigma)$, as well as $\deg_{{\mathrm{d}_{t}}}x(\sigma)\leq\deg_{{\mathrm{d}_{t}}}y(\sigma)$.
+Then there exist $q^{\star}(\sigma,{\mathrm{d}_{t}}),r^{\star}(\sigma,{\mathrm{d}_{t}})\in\mathcal{R}_{\mathrm{e}}$ such that
+$x(\sigma,{\mathrm{d}_{t}})=q^{\star}(\sigma,{\mathrm{d}_{t}})y(\sigma,{\mathrm{d}_{t}})+r^{\star}(\sigma,s)$ (29)
+with $\deg_{\sigma}r^{\star}(\sigma,{\mathrm{d}_{t}})=\deg_{\sigma}y(\sigma,{\mathrm{d}_{t}})$, $\deg^{+}_{\sigma}r^{\star}(\sigma,{\mathrm{d}_{t}})=\deg^{+}_{\sigma}y(\sigma,{\mathrm{d}_{t}})$ and $\deg^{-}_{\sigma}r^{\star}(\sigma,{\mathrm{d}_{t}})=\deg^{-}_{\sigma}y(\sigma,{\mathrm{d}_{t}})$.
 Proof:
 This proof is heavily inspired by [9, 10] and only adapted to the current setup.
-First, we obtain $q(\sigma,\partial_{t}),r(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{r}}$ by means of Lemma 2 and extract its normalised common denominator $d(\partial_{t})\in\mathbb{R}[\partial_{t}]$ such that $q(\sigma,\partial_{t})=\frac{\tilde{q}(\sigma,\partial_{t})}{d(\partial_{t})}$ with $\tilde{q}(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{p}}$.
-Next, the factorisation $d(\partial_{t})=\prod_{k=1}^{D}(\partial_{t}-s_{k})$ with the roots $s_{k}$ and $D=\deg d(\partial_{t})$ are used together with $\hat{q}_{0}(s)=\mathcal{L}\{\tilde{q}(\sigma,\partial_{t})\}$ in the iteration
+First, we obtain $q(\sigma,{\mathrm{d}_{t}}),r(\sigma,{\mathrm{d}_{t}})\in\mathcal{R}_{\mathrm{r}}$ by means of Lemma 2 and extract its normalised common denominator $d({\mathrm{d}_{t}})\in\mathbb{R}[{\mathrm{d}_{t}}]$ such that $q(\sigma,{\mathrm{d}_{t}})=\frac{\tilde{q}(\sigma,{\mathrm{d}_{t}})}{d({\mathrm{d}_{t}})}$ with $\tilde{q}(\sigma,{\mathrm{d}_{t}})\in\mathcal{R}_{\mathrm{p}}$.
+Next, the factorisation $d({\mathrm{d}_{t}})=\prod_{k=1}^{D}({\mathrm{d}_{t}}-s_{k})$ with the roots $s_{k}$ and $D=\deg d({\mathrm{d}_{t}})$ are used together with $\hat{q}_{0}(s)=\mathcal{L}\{\tilde{q}(\sigma,{\mathrm{d}_{t}})\}$ in the iteration
 $\hat{q}_{k}(s)=\frac{\hat{q}_{k-1}(s)-\hat{q}_{k-1}(s_{k})}{s-s_{k}},$ (30)
-which introduces a nominator zero for every pole from $q(\sigma,\partial_{t})$ it adds.
-Thus, $q_{D}(s)$ obtained after $D$ iterations is an entire function on $\mathbb{C}$ and therefore $q^{\star}(\sigma,\partial_{t})=\mathcal{L}^{-1}\{q_{D}(s)\}\in\mathcal{R}_{\mathrm{w}}$ by [11, Thm.
+which introduces a numerator zero for every pole from $q(\sigma,{\mathrm{d}_{t}})$ it adds.
+Thus, $q_{D}(s)$ obtained after $D$ iterations is an entire function on $\mathbb{C}$ and therefore $q^{\star}(\sigma,{\mathrm{d}_{t}})=\mathcal{L}^{-1}\{q_{D}(s)\}\in\mathcal{R}_{\mathrm{e}}$ by [21, Thm.
 7.3.1].
-Now, due to the structure of (30) we can justify the separation $q^{\star}(\sigma,\partial_{t})=q(\sigma,\partial_{t})+p(\partial_{t})$ with $p(\partial_{t})\in\mathbb{R}(\partial_{t})$.
-Therefore, we obtain the degree conditions from Lemma 2 with equal signs for the new remainder $r^{\star}(\sigma,\partial_{t})$.
+Now, due to the structure of (30) we can justify the separation $q^{\star}(\sigma,{\mathrm{d}_{t}})=q(\sigma,{\mathrm{d}_{t}})+p({\mathrm{d}_{t}})$ with $p({\mathrm{d}_{t}})\in\mathbb{R}({\mathrm{d}_{t}})$.
+Therefore, we obtain the degree conditions from Lemma 2 with equal signs for the new remainder $r^{\star}(\sigma,{\mathrm{d}_{t}})$.
 ∎
-IV-D4 Algorithm
-Equipped with a suitable way to reduce the degrees of its elements by means of the GPLD from Lemma 3, we now reduce the degrees in the entire matrix $\mathbf{H}(\sigma,\partial_{t})$ by applying a series of row transformations.
-Algorithm 1 Shift reduction in $\mathbf{H}(\sigma,\partial_{t})$
+IV-D5 Algorithm
+Equipped with a suitable way to reduce the degrees of its elements by means of the QPLD from Lemma 3, we now reduce the degrees in the entire matrix $\mathbf{G}(\sigma,{\mathrm{d}_{t}})$ by applying a series of row transformations.
+Algorithm 1 Shift reduction in $\mathbf{G}(\sigma,{\mathrm{d}_{t}})$
 alg]alg:shift_red
-1: $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]function compute_transform($\mathbf{X},j$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
+1: $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]function build_transform($\mathbf{X},j$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
 2:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\mathbf{L}_{c}\leftarrow\mathbf{I}_{{n^{-}}}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
@@
 4:   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] // From Lemma 3: $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
-5:   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $q^{\star}_{i,j}\leftarrow$ GPLD($x=\mathbf{X}[i,j],y=\mathbf{X}[j,j]$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
+5:   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $q^{\star}_{i,j}\leftarrow$ qpld($x=\mathbf{X}[i,j],y=\mathbf{X}[j,j]$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
 6:   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\mathbf{L}_{c}[i,j]\leftarrow-q^{\star}_{i,j}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
@@
 8:
-9: $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]procedure reduce_shifts($\mathbf{H}$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
-10:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\bar{}\mathbf{H}\leftarrow\mathbf{H}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
-11:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]for $i\in\{2,\dotsc,{n^{-}}\}$ do $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
-12:   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]for $j\in\{1,\dotsc,{n^{-}}-1\}$ do $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
-13:     $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\mathbf{L}_{c}\leftarrow$ Compute transform($\bar{}\mathbf{H},j$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
-14:     $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\bar{}\mathbf{H}\leftarrow\mathbf{L}_{c}\mathbf{H}$, $\mathbf{L}\leftarrow\mathbf{L}_{c}\mathbf{L}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
-15:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] return $\bar{}\mathbf{H}$$\mathbf{L}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
+9: $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]procedure reduce_shifts($\mathbf{G}$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
+10:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\bar{}\mathbf{G}\leftarrow\mathbf{G},\;\mathbf{L}\leftarrow\mathbf{I}_{{n^{-}}}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
+11:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]while $\sum_{i=0}^{{n^{-}}}\operatorname{coldeg}_{i}\mathbf{\bar{G}}>\tau_{\Sigma}$ do $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
+12:   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]if $\deg_{\sigma}\bar{g}_{ii}\neq\deg_{\sigma}\bar{g}_{jj}\;\forall i,j\in\{1,\dotsc,{n^{-}}\}$ then $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
+13:     $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]for $j\in\{1,\dotsc,{n^{-}}-1\}$ do $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
+14:      $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\mathbf{L}_{c}\leftarrow$ build_transform($\bar{}\mathbf{G},j$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
+15:      $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\bar{}\mathbf{G}\leftarrow\mathbf{L}_{c}\mathbf{G}$, $\mathbf{L}\leftarrow\mathbf{L}_{c}\mathbf{L}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
+16:   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]else $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
+17:     $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\mathbf{L}_{c}\leftarrow$ mqpld($\mathbf{\bar{G}}$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
+18:     $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\bar{}\mathbf{G}\leftarrow\mathbf{L}_{c}\mathbf{G}$, $\mathbf{L}\leftarrow\mathbf{L}_{c}\mathbf{L}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
+19:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] return $\bar{}\mathbf{G}$$\mathbf{L}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
 Lemma 4
-The matrix $\bar{}\mathbf{H}(\sigma,\partial_{t})=\mathbf{L}(\sigma,\partial_{t})\mathbf{H}(\sigma,\partial_{t})$ obtained by LABEL:alg:shift_red satisfies the requirements (24).
+The matrix $\bar{}\mathbf{G}(\sigma,{\mathrm{d}_{t}})=\mathbf{L}(\sigma,{\mathrm{d}_{t}})\mathbf{G}(\sigma,{\mathrm{d}_{t}})$ obtained by LABEL:alg:shift_red satisfies the requirements (23).
+Proof:
+In the while loop, for the case of unique pivot degrees, the degrees in each column are corrected separately from left to right.
+Therein, each correction step can introduce new shifts to the rows below its pivot element.
+In the case of coinciding degrees of pivot elements, a matrix-valued quasipolynomial long division (MQPLD) is performed which reduces the degrees in several columns at once but is not discussed here due to space limitations.
+Then the loop repeats, thus, effectively reducing the delays in each column by the (nonzero) degree differences of the pivot elements until the desired state dimension is reached.
+∎
+IV-E Deriving a hyperbolic MIMO controller form
+Applying LABEL:alg:shift_red to $\mathbf{G}(\sigma,{\mathrm{d}_{t}})$ from (20) we obtain
+$\mathbf{L}^{-1}(\sigma,{\mathrm{d}_{t}})\bm{u}(t)=\bar{}\mathbf{G}(\sigma,{\mathrm{d}_{t}})\bm{y}(t)$ (31)
+with $\hat{\mathbf{G}}=\operatorname{lccm}_{{\mathrm{d}_{t}},\sigma}\bar{}\mathbf{G}(\sigma,{\mathrm{d}_{t}})$ having full rank.
+Thus, by means of the separation $\mathbf{\bar{G}}(\sigma,{\mathrm{d}_{t}})=\mathbf{\hat{G}}\mathbf{K}(\sigma,{\mathrm{d}_{t}})+\mathbf{\check{G}}(\sigma,{\mathrm{d}_{t}})$ where $\mathbf{K}(\sigma,{\mathrm{d}_{t}})=\operatorname{diag}({\mathrm{d}_{t}}^{\nu_{1}}\sigma^{\rho_{1}},\dotsc,{\mathrm{d}_{t}}^{\nu_{m}}\sigma^{\rho_{m}})$ with the highest derivative and prediction orders $\nu_{i}=\deg_{{\mathrm{d}_{t}}}\bar{g}_{ii}(\sigma,{\mathrm{d}_{t}})$ and $\rho_{i}=\deg^{+}_{\sigma}\bar{g}_{ii}(\sigma,{\mathrm{d}_{t}})$ we obtain the explicit FDE system
+$\mathbf{K}(\sigma,{\mathrm{d}_{t}})\bm{y}(t)\!=\!\hat{\mathbf{G}}^{-1}\!\!\left(\mathbf{L}^{-1}(\sigma,{\mathrm{d}_{t}})\bm{u}(t)\!-\tilde{\mathbf{G}}(\sigma,{\mathrm{d}_{t}})\bm{y}(t)\!\right)$ (32)
+which reads
+$y_{i}^{(\nu_{i})}(t+\rho_{i})=-\sum_{j=1}^{m}\left((\tilde{a}^{\star}_{ij}y_{j})(t)-(\tilde{b}^{\star}_{ij}u_{j})(t)\right)$ (33)
+for $i=\{1,\dotsc,m\}$ in the analytic setting.
+Herein, the operators $\tilde{a}^{\star}_{ij}$ and $\tilde{b}^{\star}_{ij}$ by construction have compact support, which is why their application yields definite integrals whose limits depend on the degrees of $\tilde{\mathbf{G}}(\sigma,{\mathrm{d}_{t}})$ and $\mathbf{L}^{-1}(\sigma,{\mathrm{d}_{t}})$ w.r.t.
+$\sigma$, respectively.
+Thus, (33) states a well-posed initial value problem if the components $y_{i}(t+\tau)$ of the flat output are given for $\tau\in[\delta_{i},\rho_{i}]$ with $\delta_{i}=\deg^{-}_{\sigma}\bar{g}_{ii}(\sigma,{\mathrm{d}_{t}})$ and the input $\bm{u}(t)$ is known for $t\in\mathbb{R}$.
+Hence, we introduce the HCF state
+$\eta^{\mathrm{O}}_{i,j}(t)$ $\coloneq y_{i}^{(j-1)}(t),\;j\in\{1,\dotsc,\nu_{i}\}$ (34a)
+$\eta^{\mathrm{P}}_{i}(\tau,t)$ $\coloneq y_{i}^{(\nu_{i})}(t+\delta_{i}+\tau),\;\tau\in[0,\hat{}\tau_{i}]$ (34b)
+with $\hat{}\tau_{i}=\rho_{i}-\delta_{i}$ for $i\in\{1,\dotsc,m\}$ which is equivalent to the flat output on these intervals.
+Substituting [^5] Some expressions in $\bm{y}(t)$ may remain as e.g.
+the values $y_{i}^{(d)}(t+\tau)$ are not part of the HCF state $\bm{\eta}(\cdot,t)$ for $d<\nu_{i}$, cf.
 (12).
-Proof:
-Starting in the second row, the degrees in each column are corrected from left to right, where each of these correction steps will introduce new undesired shifts to its left in the rows below its diagonal element.
-Therefore, the procedure then continues to the next row, until no more new shifts are introduced as the last row is reached.
-∎
-IV-E Hyperbolic MIMO controller form
-Applying LABEL:alg:shift_red to $\mathbf{H}(\sigma,\partial_{t})$, we obtain
-$\mathbf{L}^{-1}(\sigma,\partial_{t})\bm{u}(\partial_{t})=\bar{}\mathbf{H}(\sigma,\partial_{t})\bm{y}(\partial_{t})$ (31)
-with $\hat{\mathbf{H}}=\operatorname{lccm}_{s,\sigma}\bar{}\mathbf{H}(\sigma,\partial_{t})$ having full rank.
-Thus, by means of the separation $\bar{}\mathbf{H}(\sigma,\partial_{t})=\hat{}\mathbf{H}\mathbf{K}(\sigma,\partial_{t})+\tilde{}\mathbf{H}(\sigma,\partial_{t})$ where $\mathbf{K}(\sigma,\partial_{t})=\operatorname{diag}(\partial_{t}^{\nu_{1}}\sigma^{\hat{\tau}_{1}},\dotsc,\partial_{t}^{\nu_{m}}\sigma^{\hat{\tau}_{m}})$ with the highest derivative and prediction orders $\nu_{i}=\deg_{\partial_{t}}\bar{h}_{i,i}(\sigma,\partial_{t})$ and $\hat{\tau}_{i}=\deg^{+}_{\sigma}\bar{h}_{i,i}(\sigma,\partial_{t})$ we obtain the explicit DDE system
-$\mathbf{K}(\sigma,\partial_{t})\bm{y}(\partial_{t})=\hat{\mathbf{H}}^{-1}\!\left(\mathbf{L}^{-1}(\sigma,\partial_{t})\bm{u}(\partial_{t})\!-\tilde{\mathbf{H}}(\sigma,\partial_{t})\bm{y}(\partial_{t})\right)$ (32)
-which reads
-$y_{i}^{(\nu_{i})}(t+\hat{\tau}_{i})=-\mathcal{A}_{i}(t)*\bm{y}(t)+\mathcal{B}_{i}(t)*\bm{u}(t)$ (33)
-for $i=\{1,\dotsc,m\}$ in the analytic setting.
-Herein, the elements of $\mathcal{A}_{i},\mathcal{B}_{i}$ by construction have compact support, which is why the convolutions will yield definite integrals whose intervals depend on the degrees of $\tilde{\mathbf{H}}(\sigma,\partial_{t})$ and $\mathbf{L}^{-1}(\sigma,\partial_{t})$ w.r.t $\sigma$.
-Thus, (33) states a well posed initial value problem if the components $y_{i}(t+\tau_{i})$ of the flat output are known for $\tau_{i}\in[\check{\tau}_{i},\hat{\tau}_{i}]$ with $\check{\tau}_{i}=\deg^{-}_{\sigma}\bar{h}_{i,i}(\sigma,\partial_{t})$ and $\bm{u}$ is known on $\mathbb{R}$.
-Hence, we introduce the HCF state
-$\eta^{\mathrm{O}}_{i,j}(t)$ $\coloneq y_{i}^{(j-1)}(t-\check{\tau}_{i}),\;j\in\{1,\dotsc,\nu_{i}\}$ (34a)
-$\eta^{\mathrm{P}}_{i}(\tau,t)$ $\coloneq y_{i}^{(\nu_{i})}(t+\tau),\;\tau\in[\check{\tau}_{i},\hat{\tau}_{i}]$ (34b)
-for $i\in\{1,\dotsc,m\}$ which is equivalent to the flat output on these intervals.
-Substituting [^5] Some expressions in $\bm{y}(t)$ may remain as e.g.
-the values $y_{i}^{(d)}(t+\tau_{i})$ are not part of the HCF state $\bm{\eta}(\cdot,t)$ for $p<\nu_{i}$, cf.
-(12).
-However, these expressions can be easily eliminated by means shown e.g.
-in [6, Lemma 5].
-the definitions (34) in (33) then allows to identify the functions $a^{\star}_{i,j}(t),b^{\star}_{i,j}(t)$ as well as weights $\bm{a}_{i}$ in (12), yielding the HCF of system (15).
+However, these expressions can be easily eliminated by means shown in [22, Lem.
+5].
+the definitions (34) in (33) then allows to identify the operators $a^{\star}_{ij},b^{\star}_{ij}$ as well as weights $\bm{a}_{ij}$ in (12), yielding the HCF of system (15).
 Lemma 5
@@
 Proof:
-The proof of Lemma 2 yields $\deg_{\partial_{t}}q(\sigma,\partial_{t})<0$ due to the structure of $d(\partial_{t})$.
-Then by Lemma 3, $q^{\star}(\sigma,\partial_{t})$ has non-positive degree as do all entries in $\mathbf{L}_{i}(\sigma,\partial_{t})$ from LABEL:alg:shift_red.
-Now, as the inverse of the triangular $\mathbf{L}(\sigma,\partial_{t})$ is computed by forward substitution, all $\bar{q}_{i,j}(\sigma,\partial_{t})$ therein are linear combinations of the $q_{i,j}^{\star}(\sigma,\partial_{t})$, hence, $\deg_{\partial_{t}}\bar{q}_{i,j}(\sigma,\partial_{t})\leq 0$.
-Therefore, the convolution with $\mathcal{B}_{i}$ yields no derivatives.
+The proof of Lemma 2 yields $\deg_{{\mathrm{d}_{t}}}q(\sigma,{\mathrm{d}_{t}})<0$ due to the structure of $d({\mathrm{d}_{t}})$.
+Then by Lemma 3, $q^{\star}(\sigma,{\mathrm{d}_{t}})$ has non-positive degree w.r.t.
+$\mathrm{d}_{t}$} as do all entries in $\mathbf{L}_{i}(\sigma,{\mathrm{d}_{t}})$ from LABEL:alg:shift_red.
+Now, as the inverse of the triangular $\mathbf{L}(\sigma,{\mathrm{d}_{t}})$ is computed by forward substitution, all $\bar{q}^{\star}_{ij}(\sigma,{\mathrm{d}_{t}})$ therein are linear combinations of the $q_{ij}^{\star}(\sigma,{\mathrm{d}_{t}})$, hence, $\deg_{{\mathrm{d}_{t}}}\bar{q}^{\star}_{ij}(\sigma,{\mathrm{d}_{t}})\leq 0$.
+Therefore, the application of the $\tilde{b}^{\star}_{ij}$ yields (distributed) shifts but no derivatives.
 ∎
 V Reduction of the benchmark system
-Consider again the system from Section II-C with $\alpha=\pi$, $\beta=10$ and $m=1$ substituted into (10) yielding
-$\mathbf{H}(\sigma,\partial_{t})=\frac{1}{4}\begin{pmatrix}a_{1}(\partial_{t})\sigma^{\pi}+a_{2}(\partial_{t})\sigma^{-\pi}&-\sigma^{\pi}-\sigma^{-\pi}\\ b_{1}(\partial_{t})\sigma^{10}+b_{2}(\partial_{t})\sigma^{-10}&\sigma^{10}+\sigma^{-10}\\ \end{pmatrix}$
-with $a_{1}(\partial_{t})=b_{1}(\partial_{t})=\partial_{t}^{2}+2\partial_{t}$ and $a_{2}(\partial_{t})=b_{2}(\partial_{t})=\partial_{t}^{2}-2\partial_{t}$ for which $\operatorname{lccm}_{\sigma,\partial_{t}}\mathbf{H}(\sigma,\partial_{t})=\operatorname{lccm}_{\partial_{t},\sigma}\mathbf{H}(\sigma,\partial_{t})$ is of rank $1$ where $2$ would be needed.
-Now, regarding Section IV-D1, the rows are already correctly ordered since $\pi<10$ and also the columns do not need to be reordered since the coefficients in each column are of the same degree w.r.t.
-$\partial_{t}$.
-Continuing with Section IV-D2, the conditions (25) fail for $h_{2,1}(\sigma,\partial_{t})$ as e.g.
-$\deg_{\sigma}h_{2,1}(\sigma,\partial_{t})=20>\deg_{\sigma}h_{1,1}(\sigma,\partial_{t})=2\pi$.
+Consider again the system from Section II-C with ${l_{1}}=\pi$, ${l_{2}}=10$ and M=1{M}=1 substituted into (10) yielding
+$\mathbf{G}(\sigma,{\mathrm{d}_{t}})=\frac{1}{4}\begin{pmatrix}a_{1}({\mathrm{d}_{t}})\sigma^{\pi}+a_{2}({\mathrm{d}_{t}})\sigma^{-\pi}&-\sigma^{\pi}-\sigma^{-\pi}\\ b_{1}({\mathrm{d}_{t}})\sigma^{10}+b_{2}({\mathrm{d}_{t}})\sigma^{-10}&\sigma^{10}+\sigma^{-10}\\ \end{pmatrix}$
+with $a_{1}({\mathrm{d}_{t}})=b_{1}({\mathrm{d}_{t}})={\mathrm{d}_{t}}^{2}+2{\mathrm{d}_{t}}$ and $a_{2}({\mathrm{d}_{t}})=b_{2}({\mathrm{d}_{t}})={\mathrm{d}_{t}}^{2}-2{\mathrm{d}_{t}}$ for which $\hat{}\mathbf{G}=\operatorname{lccm}_{\sigma,{\mathrm{d}_{t}}}\mathbf{G}(\sigma,{\mathrm{d}_{t}})=\operatorname{lccm}_{{\mathrm{d}_{t}},\sigma}\mathbf{G}(\sigma,{\mathrm{d}_{t}})$ is singular as shown in (11) which is why the degree conditions (24) from Section IV-D1 fail; e.g.
+$\deg_{\sigma}g_{21}(\sigma,{\mathrm{d}_{t}})=20>\deg_{\sigma}g_{11}(\sigma,{\mathrm{d}_{t}})=2\pi$.
+Continuing with Section IV-D2, no superfluous delays are present, rendering the delay triangularisation and homogenisation obsolete.
+Furthermore, the row degrees w.r.t.
+$\sigma$ are already ordered and the columns do not need to be reordered as their entries satisfy the degree condition (24b).
 Therefore, by means of Lemma 2 the quotient
-$\begin{aligned} q_{1,2}&(\sigma,\partial_{t})\!=\!\sigma^{10-\pi}\!+\!\frac{2-\partial_{t}}{\partial_{t}+2}\sigma^{10-3\pi}\!-\!\frac{\partial_{t}+2}{\partial_{t}-2}\sigma^{-10+3\pi}\!+\!\sigma^{-10+\pi}\end{aligned}$
+$\begin{aligned} q_{21}&(\sigma,{\mathrm{d}_{t}})\!=\!\sigma^{10-\pi}\!+\!\frac{2-{\mathrm{d}_{t}}}{{\mathrm{d}_{t}}+2}\sigma^{10-3\pi}\!-\!\frac{{\mathrm{d}_{t}}+2}{{\mathrm{d}_{t}}-2}\sigma^{-10+3\pi}\!+\!\sigma^{-10+\pi}\end{aligned}$
 is obtained, where one iteration was needed for each case in (28).
-Next, Lemma 3 yields the correction term $p(\partial_{t})=\frac{16e^{-20+6\pi}}{\left(\partial_{t}^{2}-4\right)}$ and thus $q_{2,1}^{\star}(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{w}}$.
+Next, Lemma 3 yields the correction term $p({\mathrm{d}_{t}})=\frac{16e^{-20+6\pi}}{\left({\mathrm{d}_{t}}^{2}-4\right)}$ such that $q_{21}^{\star}(\sigma,{\mathrm{d}_{t}})\in\mathcal{R}_{\mathrm{e}}$.
 Using the latter in (26) then yields the reduced element
-$\begin{aligned} \bar{h}_{1,2}(\sigma,\partial_{t})=\!-\frac{16\partial_{t}}{\left(\partial_{t}-2\right)e^{20-6\pi}}\sigma^{\pi}\!+\frac{\partial_{t}\left(\partial_{t}^{2}+4\partial_{t}+4\right)}{\partial_{t}-2}\sigma^{-10+4\pi}\\ \!+\frac{\partial_{t}\left(\partial_{t}^{2}-4\partial_{t}+4\right)}{\partial_{t}+2}\sigma^{10-4\pi}\!-\frac{16\partial_{t}}{\left(\partial_{t}+2\right)e^{20-6\pi}}\sigma^{-\pi}.\end{aligned}$
-As the requirements for the second column are already satisfied, this concludes the reduction and we obtain $\hat{}\mathbf{H}=\operatorname{lccm}_{\partial_{t},\sigma}\bar{}\mathbf{H}(\sigma,\partial_{t})=\operatorname{diag}(1,-2)$.
-Python code showing the complete computation of the HCF will be made available at a later time.
+$\begin{aligned} \bar{g}_{21}(\sigma,{\mathrm{d}_{t}})=\!-\frac{16{\mathrm{d}_{t}}}{\left({\mathrm{d}_{t}}-2\right)e^{20-6\pi}}\sigma^{\pi}\!+\frac{{\mathrm{d}_{t}}\left({\mathrm{d}_{t}}^{2}+4{\mathrm{d}_{t}}+4\right)}{{\mathrm{d}_{t}}-2}\sigma^{-10+4\pi}\\ \!+\frac{{\mathrm{d}_{t}}\left({\mathrm{d}_{t}}^{2}-4{\mathrm{d}_{t}}+4\right)}{{\mathrm{d}_{t}}+2}\sigma^{10-4\pi}\!-\frac{16{\mathrm{d}_{t}}}{\left({\mathrm{d}_{t}}+2\right)e^{20-6\pi}}\sigma^{-\pi}.\end{aligned}$
+with $\deg_{\sigma}\bar{g}_{21}(\sigma,{\mathrm{d}_{t}})=\deg_{\sigma}g_{11}(\sigma,{\mathrm{d}_{t}})$.
+As the requirements for the second column are already satisfied, this concludes the reduction and we obtain $\mathbf{\hat{G}}=\operatorname{lccm}_{{\mathrm{d}_{t}},\sigma}\mathbf{\bar{G}}(\sigma,{\mathrm{d}_{t}})=\operatorname{diag}(1,-2)$.
+[^6] The complete computation of the HCF for the example with annotated steps can be found at https://github.com/umit-iace/iace-pub-2026-ecc.
 VI Concluding remarks
@@
 Furthermore, the obtained HCF contains input delays, which will result in a dynamic extension, thus further analysis is required to obtain a form without delays as this would enable the application of quasi-static state feedback.
-Finally, the system class (15) is lacking in-domain couplings, as these were recently addressed in [13] for the MIMO case without boundary dynamics, the next logical step is to combine the two approaches.
+Finally, the system class (15) is lacking in-domain couplings.
+As this class was recently addressed in [23] without boundary dynamics, the next logical step is to combine the two approaches.